### v2

# A Survey on Speech Large Language Models

Abstract
Large Language Models (LLMs) exhibit strong contextual understanding and remarkable multi-task performance. Therefore, researchers have been seeking to integrate LLMs in the broad sense of Spoken Language Understanding (SLU) field. Different from the traditional method of cascading LLMs to process text generated by Automatic Speech Recognition(ASR), new efforts have focused on designing architectures centered around Audio Feature Extraction - Multimodal Information Fusion - LLM Inference (Speech LLMs). This approach enables richer audio feature extraction while simultaneously facilitating end-to-end fusion of audio and text modalities, thereby achieving deeper understanding and reasoning from audio data. This paper elucidates the development of Speech LLMs, offering an in-depth analysis of system architectures and training strategies. Through extensive research and a series of targeted experiments, the paper assesses Speech LLMs’ advancements in Rich Audio Transcription and its potential for Cross-task Integration within the SLU field. Additionally, it indicates key challenges uncovered through experimentation, such as the Dormancy of LLMs under certain conditions. The paper further delves into the training strategies for Speech LLMs, proposing potential solutions based on these findings, and offering valuable insights and references for future research in this domain, as well as LLM applications in multimodal contexts.
1 Introduction
The advancements in large language models within natural language processing have been particularly significant, especially in contextual understanding and multimodal technology. Models like GPT-4 [1], trained on extensive corpora, are now capable of processing complex contexts with exceptional accuracy. These models, with their vast parameters, excel in parsing contextually appropriate sentences and maintaining coherence over multiple conversational turns, which is crucial for tasks like dialogue systems, automatic summarization, and machine translation. By leveraging vast contextual information, these models manage long-text processing with unprecedented proficiency in contextual reasoning.
In addition to the success in text-based tasks, large language models have also achieved remarkable success in multimodal tasks [2], particularly in integrating and processing inputs from diverse modalities like vision, speech, and text. Their ability to semantically link these different types of data has led to significant breakthroughs in areas such as visual question answering and image generation, as exemplified by models like CLIP [3] and DALL-E [4]. This capability not only enhances performance in visual tasks by effectively fusing information from different modalities, but also demonstrates the models’ potential to expand into and improve upon other modalities. The way these models manage complex multimodal information processing highlights their adaptability and lays the groundwork for further innovations in multimodal understanding and integration.
The typical Spoken Language Understanding (SLU) involves interpreting spoken language to extract meaning, intent, and relevant information beyond simple transcription [5, 6]. In order to more comprehensively and coherently demonstrate the performance of Speech LLMs, in this article, we also include basic speech-to-text tasks, such as automatic speech recognition (ASR), within its scope. Related works have achieved remarkable performance across various domains, including voice-activated assistants, automated transcription, and real-time speech translation. Modern systems are adept at handling diverse accents and languages, improving efficiency and accuracy in workflows like medical transcription and customer service automation. However, challenges remain, particularly in long-form speech recognition and hotword/keyword recognition. Long-form recognition struggles with maintaining context over extended periods and handling speaker variability, often resulting in accuracy degradation and latency issues in real-time applications. Hotword detection, critical for wake word recognition, faces difficulties in noisy environments and balancing sensitivity and specificity, especially when hotwords are contextually similar to other phrases. The integration of hotword recognition into continuous speech processing without disrupting flow or increasing computational overhead remains an ongoing challenge.
In this paper, our contributions can be summarized in the following three points:
  1. 1.
    The first in-depth and comprehensive survey analyzing Speech LLMs in the SLU domain. We will conduct an in-depth exploration of the development of Speech LLMs, explaining the paradigm of model architecture, and perform a comparative analysis with traditional speech models in SLU tasks.
  2. 2.
    The first comparative analysis of training methods for aligning speech and text modalities in Speech LLMs architecture, with an emphasis on the potential development of Reinforcement Learning strategies such as DPO and PPO.
  3. 3.
    We are the first to propose and deeply analyze the LLM’s Dormancy when the LLM is applied in the speech domain, and to provide possible viable conclusions regarding its causes.
2 Recent Advances in Speech LLMs
2.1 Evolution of Speech LLMs Architectures
Considering the exceptional performance of large language models (LLMs) in contextual understanding, reasoning, and multimodal information processing, and recognizing the unresolved challenges in the field of SLU, such as long-form speech understanding and hotword recognition, researchers have begun to explore the integration of LLMs into SLU.
2.1.1 Transformer Integration into Traditional Speech Models
One of the key milestones in the integration of LLMs into SLU was the introduction of the Transformer architecture. Dong et al. (2018) first applied the Transformer to speech recognition, proposing an end-to-end speech recognition system based on the Transformer [7]. This attempt laid the foundation for subsequent research. Following this, in 2020, Gulati et al. (2020) further advanced this approach by introducing the Conformer architecture, which combines the local feature extraction capabilities of Convolutional Neural Networks (CNNs) with the Transformer, making the Transformer more robust in processing speech signals [8].
In 2021, Hsu et al. proposed HuBERT (Hidden-Unit BERT, which represented a major breakthrough in utilizing LLMs to process audio features [9]. HuBERT employs self-supervised learning on a large corpus of unlabeled speech data to learn effective hidden unit representations. In 2022, Radford et al. introduced the Whisper model, which further integrated multilingual and multitask capabilities into a single model, demonstrating excellent transferability across different languages and tasks [3].
Researchers, unsatisfied with simply using speech as input and text as output, began exploring end-to-end speech-text multimodal inputs and outputs. In 2022, Junyi Ao et al. introduced SpeechT5, pioneering the investigation of a unified encoder-decoder framework for various spoken language processing tasks [10].
2.1.2 Direct Audio Processing with Large Language Models
Fig. 1: The evolution of Speech LLMs from the Pre-LLM Stage with separate acoustic and language models, to the Early LLM Stage with transformer integration, and the Current Speech LLM Stage, featuring direct audio processing by large language models. The current stage includes two branches: Discrete Sequence Modeling (e.g., VALL-E, SpeechGPT) and Continuous Sequence Modeling (e.g., Pengi, SALMONN).
Besides directly incorporating the architecture of LLMs like transformer into speech recognition tasks, another significant integration approach involves using LLM as a whole to process audio features extracted by traditional speech recognition tools. This approach leverages the powerful contextual understanding and reasoning abilities of LLMs to improve the accuracy and robustness of speech recognition and deeper speech understanding, and it has gradually evolved into the main trend of using large language models in the field of speech recognition.In other words, this refers to the narrower definition of Speech LLMs .
The earliest attempts to align speech and text modalities by passing extracted features from speech to downstream language models can be traced back to 2020. Qi Liu et al. were the first to explore this approach, separating the mapping of speech feature information from the speech feature extraction process into two independent modules and introducing the LSTM model into the process of mapping speech feature information to text [11]. This approach differs from traditional CTC-based end-to-end systems by isolating the acoustic feature-to-text mapping as a standalone unit. This separation laid the groundwork for the subsequent integration of large language models into this area, making it possible to embed LLMs more effectively in speech recognition tasks.
Currently, Speech LLMs can mainly be divided into two categories 4. The first is Discrete Sequence Modeling, represented by approaches that condense audio feature information into discrete tokens, which are then passed to the LLM for processing. In early 2023, Microsoft released Vall-E, a typical study utilizing LLMs to process audio features. Vall-E introduces the Transformer architecture, combining audio features with the capabilities of language models to achieve more natural text-to-speech generation [12]. Following this, the emergence of SpeechGPT marked a deep integration of speech models and LLMs, capable not only of processing audio but also interacting through natural language, bringing a new interactive paradigm to the field of speech recognition [13]. In mid-2023, Google launched AudioPaLM, which further expanded the capabilities of speech recognition into multimodal processing, enhancing performance in multimodal tasks [14].
The second is Continuous Sequence Modeling, where audio feature information is projected into continuous input embedding vectors and then transmitted to the LLM for processing. In 2023, Deshmukh et al. introduced Pengi, which projects speech modality information into the text space of large models without altering any parameters in LLM such as embedding vector spaces [15]. Subsequently, continuous audio feature information began to be used to convey richer representations to the LLM [16, 17, 18], resulting in significant performance improvements across a range of tasks in the SLU domain. In 2024, Ma et al. further proposed SLAM-LLM, which involves the addition of a linear projector, allowing the task to be greatly completed by training only the projection layer [19].
Once researchers achieved significant improvements in recognition accuracy compared to traditional speech recognition technologies, they began exploring and attempting to address the issues inherent in these conventional methods. In 2024, Fathullah et al. introduced the Conformer mechanism, achieving advancements in handling long speech sequences [20]. In the same year, SpeechX achieved breakthroughs in multilingual, multitask speech recognition, enabling seamless switching between languages and supporting challenges like long-form speech understanding and hotword recognition [21].
2.2 Advancements of Speech LLMs in Key Tasks and Challenges
With the continued advancement and application of large language models (LLMs) in the field of speech processing, the Speech LLM paradigm has demonstrated significant success across various tasks, often surpassing the performance of traditional models in many areas. Our primary focus is on traditional key tasks within the domain of speech understanding, which include automatic speech recognition (ASR), speech translation, and speaker identification, among others. A more detailed comparison and analysis will be presented in  Section 6. At the same time, we also focus on the exploration and breakthroughs of Speech LLM in challenging tasks such as long-form speech recognition, as well as keyword and hotword detection. Of course, the powerful contextual reasoning capabilities of large language models provide an alternative solution for more complex system tasks. In this regard, we have also explored areas such as real-time multimodal interaction. This section investigates both the performance of Speech LLMs on these standard tasks and how they address long-standing challenges in the field.
2.2.1 Improvement in Traditional Tasks in Spoken Language Understanding
Traditional tasks in the field of speech understanding include automatic speech recognition, speaker identification, and speech translation. The Speech LLM paradigm has demonstrated impressive performance in all these areas. Here, we will firstly take ASR as an example to briefly examine and understand the progress Speech LLM has made in this field.
Automatic Speech Recognition (ASR) is the technology that enables machines to convert spoken language into text. Modern ASR systems, enhanced by LLMs, aim to achieve higher accuracy, better noise resilience, and greater adaptability to diverse accents and dialects compared to their predecessors. This task is foundational for voice-controlled applications, interactive voice response systems, and automated transcription services. One of the key metrics for evaluating speech recognition systems is the Word Error Rate (WER).
Traditional models, such as those based on LSTM or GRU, have shown competitive performance on various benchmark datasets [22]. However, the introduction of LLMs has significantly improved these results [23, 24].
Moreover, in multilingual speech recognition, LLMs have demonstrated superior performance across various languages. On the Multilingual LibriSpeech (MLS) dataset, Conformer-based models have achieved significant reductions in WER across multiple languages [25]. Notably, in low-resource languages such as French and German, great WER reductions have been reported [23].
Speech translation is the process of converting spoken language from one language into written or spoken text in another language. It typically involves two key steps [26]: automatic speech recognition (ASR), which transcribes spoken words into text, and machine translation (MT), which translates the transcribed text into the target language. This technology is used in real-time applications such as multilingual meetings, conferences, and live broadcasts, enabling seamless cross-lingual communication. With the successful application of large language models in the field of machine translation [27], the speech translation domain has also begun to gradually incorporate large language models, aiming to achieve similar significant results. Advances in large language models have significantly improved the accuracy and fluency of speech translation systems. In the field of speech translation, the integration of large language models has significantly enhanced performance. Not only has it improved the accuracy of speech translation tasks [28], but it has also broadened the range of supported languages [29], making these models an invaluable addition to the field. The incorporation of large language models brings a complementary advantage, enriching both precision and diversity in multilingual speech translation.
Additionally, LLMs have excelled in multitask learning scenarios. For example, the Qwen-Audio model has shown impressive performance in tasks that combine speech-to-text with other modalities, such as sentiment analysis and speaker identification, reducing WER by 10% and improving sentiment recognition accuracy by 8% compared to single-task models [17].
For a detailed comparison and analysis of Speech LLMs’ performance in traditional speech tasks, please refer to Section 6.
2.2.2 Long-Form Speech Understanding
Traditional speech recognition systems have struggled with long-form speech understanding due to context loss over extended periods. This issue is particularly pronounced in audio segments longer than one minute, where traditional models often exhibit a sharp increase in WER. In contrast, LLM-based systems have significantly mitigated this problem. The Whisper model, for instance, maintains contextual consistency across long-form audio, demonstrating an 18% reduction in WER on audio segments exceeding five minutes compared to traditional models. This enhancement is critical for applications such as the transcription of speeches and lectures.
Additionally, models like UniAudio and Pengi have shown remarkable performance in maintaining low WERs across extended speech segments by integrating advanced contextual understanding [30, 31]. These models have been particularly effective in domains requiring detailed, accurate transcriptions of prolonged speech events, where maintaining context over extended durations is crucial.
2.2.3 Hotword Recognition
Hotword recognition has been another challenging area for traditional speech recognition systems, especially in noisy environments. Traditional models often fail to accurately detect and differentiate hotwords within long audio streams. The GenTranslate model, introduced in late 2023, has made substantial improvements in this area by leveraging the contextual understanding capabilities of LLMs [28]. In a test set containing multiple domain-specific hotwords, GenTranslate achieved a 22% improvement in hotword recognition accuracy compared to traditional models. Moreover, it maintained high robustness in noisy conditions, marking a significant advancement over prior approaches.
Further examples of LLM-based models outperforming traditional systems include Mala-ASR and Whisper, which not only improves hotword recognition accuracy [32] but also adapts dynamically to new hotwords in real-time, a feature that traditional models lack. This capability is particularly valuable in dynamic environments like live broadcasts or interactive voice response (IVR) systems where new terms may frequently be introduced.
2.2.4 Real-time Multimodal Interaction
The integration of LLMs into speech recognition has expanded the scope of tasks beyond traditional speech-to-text, enabling real-time multimodal interaction. Models like VoxtLM and LauraGPT facilitate seamless integration of speech with visual and textual inputs [33, 34], providing coherent and accurate multimodal outputs. This capability is particularly valuable in applications such as live transcription and synchronized translation during presentations, where both speech and visual context need to be processed simultaneously.
Moreover, LLM-based systems have introduced new functionalities, such as the generation of descriptive text, summaries, and even translations based on audio input. For instance, ViOLA can generate coherent summaries and cross-language translations with high fluency and accuracy, outperforming traditional models in both speed and quality of output. This represents a significant advancement in how speech recognition systems can interact with and interpret complex multimodal data streams [35].In addition, the integration of large language models with acoustics has demonstrated significant development potential across various specialized fields [36].
3 Model Paradigm of Speech LLMs
As a model architecture that integrates both audio and text modalities, Speech LLMs exhibits significant differences in input-output processing and model structure compared to traditional speech models. Notably, due to the significant commonality and prevailing trends, we choose to focus on the narrower definition of Speech LLMs 1, which involves the integration of LLMs as a whole into SLU tasks.
3.1 Speech LLM Input-Output Modes
Fig. 2: Overview of Speech LLM architectures corresponding to three different input-output modalities: Speech-to-Text (S2T), Speech&Text-to-Text (ST2T), and Speech&Text-to-Speech&Text (ST2ST).
Generally, the input and output modes of Speech LLMs can be categorized into three main types:
  1. 1.
    Speech to Text (S2T)
    This mode is the most straightforward and aligns closely with traditional speech recognition. The model’s encoder extracts features from the audio input, and the large language model then generates text as output. While S2T models are simple to implement, they are limited in the range of tasks they can handle since they only support speech-to-text conversion.
  2. 2.
    Speech&Text to Text (ST2T)
    The ST2T mode, the most widely used in modern speech recognition, incorporates both audio and text inputs. Typically, the text input serves as an instruction or task prompt. The model processes both modalities, fusing the audio and text information to produce a final text output. This approach enables multi-task capabilities and allows models to fully leverage the strengths of large language models, handling a wider variety of tasks than the S2T mode.
  3. 3.
    Speech&Text to Speech&Text (ST2ST)
    In this more advanced mode, a vocoder is added to the output, enabling the model to generate both text and audio outputs. This expands the model’s capabilities beyond basic speech recognition, supporting tasks such as Text-to-Speech (TTS), speech translation, and voice conversion. The ST2ST mode offers greater versatility than the other two modes, making it applicable across a broader range of speech and language tasks.
3.2 Model structure
Fig. 3: Overview of key stages of the Speech LLM Workflow
Up to now, various Speech LLM architectures have been developed, all of which are structured around three fundamental stages: Modality Feature Extraction, Modality Information Fusion, and LLM Inference. As illustrated in Figure 3, despite different input and output modalities, the major stages that a model went through is consistent.
Speech LLMs process input from both speech and text modalities using separate extractors tailored to the characteristics of each modality. The raw audio input is processed by an audio encoder, which extracts relevant acoustic features, while the text input is handled by the encoder of the large language model to capture linguistic features. Once these features are extracted, they are fused through various techniques, which will be elaborated upon in the following sections. After feature fusion, the combined multimodal information is sent to the large language model for final decoding and output generation. In the following sections, we will examine three distinct model architectures, each corresponding to a different input-output modality.
(a) Direct Projection
(b) Token Mapping
(c) Combine audio and text space
Fig. 4: Illustration of three approaches to audio-text information fusion in Speech LLMs
3.2.1 S2T
S2T models (Speech-to-Text) take speech as the input and generate text as the output. These models typically incorporate an audio encoder to extract features from the speech signal, while a text encoder is not required, as text input is absent. Therefore, to optimize model size and efficiency, a decoder-only architecture is commonly employed to generate the text output. However, aligning the audio features with the input embedding space of the large language model (decoder-only architecture) requires a conversion module. This module ensures that the audio features are properly mapped, facilitating the generation of accurate text output. As discussed in the work by Tsunoo, et al. [37], the decoder-only architecture is particularly effective for streaming end-to-end speech recognition, where efficiency and real-time performance are critical.
3.2.2 ST2S
ST2S (Speech & Text to Speech) models process both speech and text inputs by employing distinct encoders for each modality during the modal information extraction phase. Specifically, an audio encoder extracts relevant features from the speech input, while a text encoder captures features from the textual input. In the subsequent modal information fusion phase, it is essential to align the speech-derived features with the textual representation space to ensure coherent integration. The unified multimodal features are then passed to the decoder of a large language model, which carries out the inference and decoding process to generate the final output.
3.2.3 ST2ST
Speech&Text-to-Speech&Text models (ST2ST) include an additional speech modality in the output, which requires a more complex decoding mechanism during the generation phase. After modal feature extraction, modal information fusion, and large language model inference, the text output can be generated by the decoder of the large language model. However, the speech output requires a vocoder to synthesize the final speech output [38, 13].
4 Multimodal Information Fusion
Having established a preliminary understanding of the overall architecture of Speech LLMs, we will now examine the model’s key components in greater detail. Beyond the well-known stages of audio feature extraction and large language model inference, the most critical issue in Speech LLM lies in the alignment between the audio modality and the text modality. Aligning these two modalities typically requires two steps, each addressing a specific problem: The first step, which we refer to as Audio Feature Post-Process, focuses on determining what specific audio modality information is needed. The second step, called Audio and Text Connection, addresses how to effectively combine these two types of information.
4.1 Audio Feature Post-Process
In this stage, researchers currently tend to directly use the final layer output of the encoder. The current mainstream approach is to extract the output of the final layer of the encoder as the primary source of audio modality information. However, there are other various alternative methods available. For instance, some approaches involve using intermediate layer outputs to capture more granular features [39]. Others apply attention mechanisms to emphasize relevant parts of the audio signal [40].
4.2 Audio and Text Connection
Once the audio modality information is obtained, it must be integrated with the textual modality information to enable the LLM to perform the final inference. Currently, the methods used can be broadly classified into two categories: transforming the audio feature into the textual modality space, or merging the audio and textual modality spaces.
4.2.1 Audio-to-Text Modality Conversion
Converting speech modality information into the text modality space is one of the most commonly used approaches today. This method fully considers that large language models (LLMs) are primarily designed for text modalities, enabling direct projection and integration of the audio modality into the text modality space. This effectively minimizes modifications to the LLM while preserving its powerful capabilities. To achieve this, a connector or projector is needed to transform or project the extracted audio modality features. This neural network structure should minimize the loss of audio feature information while ensuring smooth transfer from the audio feature space to the text feature space.
Currently, two common methods are employed (illustrated in  Figure 4(a) and Figure 4(b)):
  1. 1.
    Direct Projection: The audio feature information is directly projected into the LLM’s text feature space through a connector [19, 16]. Specifically, the tensor containing audio features is projected into an embedding vector aligned with the text modality. These audio embeddings are then concatenated with the input text’s embedding vector to create a new embedding vector that integrates both speech and text information. This combined vector is then fed into the LLM.Of course, some researchers have also implicitly handled the projection step by internalizing it within the original encoder, achieving the modality projection effect by adjusting the encoder’s parameters during training [17].
  2. 2.
    Token Mapping: The audio feature information is mapped into text tokens that the LLM can process [37]. Initially, the audio features are passed through a projector or converter to generate symbols or representations that correspond to text tokens. These representations are then combined with the tokenized input text, forming a token sequence that includes both audio and text information. This sequence is subsequently input into the LLM for processing. This approach not only preserves the integrity of the audio feature information but also ensures consistent processing within the LLM.
4.2.2 Combining Audio and Text Feature Space
Projecting the speech modality into the text modality is certainly straightforward, but it does not achieve lossless modality fusion in the true sense. Information loss and conflicts may occur during modality conversion (to be elaborated later). Therefore, researchers have proposed another modality alignment approach, which involves modifying the original input space of the large language model to integrate the audio modality [14, 13, 41]. Specifically, this method augments the token space by adding audio tokens on top of the existing text tokens, creating a new token space. These new audio token are synthesized from the audio features extracted in the previous step, thereby largely preserving the original audio information (illustrated in Figure 4(c)).
5 Training Strategies
The training of current Speech LLMs mainly includes three approaches: pretraining, supervised fine-tuning (SFT), and reinforcement learning (RL). These approaches are used in different training stages and scenarios due to their varying principles and effects.
5.1 Pretraining
Pretraining involves training models on large-scale unlabeled data to enable them to learn broad, general knowledge. Typically, in Speech LLMs, speech encoders undergo pretraining on audio-text pairs to capture audio features. Common training strategies, including self-supervised learning (SSL) [18], are widely used. However, to better integrate speech encoders with LLMs, some researchers attempt to re-pretrain speech encoders [42], which allows for more direct acquisition of encoders with audio feature extraction capabilities.
For Speech LLMs of Section 4.2.2, as mentioned previously, a thorough re-training of multimodal large models is necessary. This involves referencing the pretraining phase of large language models. By organizing mixed audio and text data, we adapt the pretraining methods of large language models to accommodate bimodal information.
5.2 Supervised Fine-Tuning (SFT)
For specific tasks in the speech domain, further fine-tuning is often required. Supervised fine-tuning is a common approach where high-quality labeled data from downstream task datasets is used to train the model, improving its performance on these tasks. In Speech LLMs, for models of structure as mentioned earlier in Section 4.2.1, supervised fine-tuning is an extremely common method. It is generally used to achieve alignment between the speech encoder and the LLM and to enhance performance on specific tasks. Common training methods include fine-tuning connectors, fine-tuning the encoder, and LLMs, such as using methods like LoRA. This mainly involves handling modality alignment and completing the model’s learning of text-token mapping. Similarly, models of structure described in  Section 4.2.2 also require supervised fine-tuning, which is essential for enhancing its performance on specific tasks.
It is worth noting that supervised fine-tuning is also frequently used as a wake-up method for large language models. When modality alignment is completed, large language models might experience a certain degree of dormancy [16]. Researchers use audio-based dialogue wake-up fine-tuning to reactivate large language models.
5.3 Reinforcement Learning (RL)
Reinforcement learning (RL) is a commonly used method in the training of large language models, especially in the field of large language model safety alignment [43]. Unlike other training strategies, reinforcement learning ensures that the LLM optimizes in the desired direction while maintaining diversity in its outputs. In order to maintain and enhance cross-task integration capabilities of large language models under the Speech LLM paradigm, reinforcement learning is an effective training strategy that is worth further exploration [44, 45]. However, it has not been widely used in Speech LLMs.
Reinforcement learning is a framework where an agent interacts with an environment to maximize cumulative rewards over time. The agent learns by receiving feedback in the form of rewards, adjusting its actions to improve future rewards. Two commonly used RL algorithms in this context are Proximal Policy Optimization (PPO) and Direct Policy Optimization (DPO).
Proximal Policy Optimization (PPO) is a policy gradient method that aims to optimize the objective function while limiting policy updates to avoid large deviations from the current policy. The objective function for PPO is defined as:
$\mathcal{L}^{\text{PPO}}(\theta)=\hat{\mathbb{E}}_{t}\left[\min\left(r_{t}(\theta)\hat{A}_{t},\text{clip}(r_{t}(\theta),1-\epsilon,1+\epsilon)\hat{A}_{t}\right)\right]$ (1)
where $r_{t}(\theta)=\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{\text{old}}}(a_{t}|s_{t})}$ is the probability ratio between the new policy and the old policy, $\hat{A}_{t}$ is the advantage function, and $\epsilon$ is a hyperparameter that controls the clipping range. The clipping mechanism ensures that policy updates remain within a trust region, preventing overly aggressive updates that could harm the model’s performance.
Direct Policy Optimization (DPO), on the other hand, focuses on directly optimizing the policy by defining a reward function specific to the task. The objective in DPO can be expressed as:
$\mathcal{L}^{\text{DPO}}(\theta)=\hat{\mathbb{E}}_{t}\left[R(s_{t},a_{t})\log\pi_{\theta}(a_{t}|s_{t})\right]$ (2)
where $R(s_{t},a_{t})$ is the task-specific reward function, and $\pi_{\theta}(a_{t}|s_{t})$ is the policy parameterized by $\theta$. DPO directly maximizes the expected reward by focusing on optimizing the policy without clipping, often used in settings where a more direct reward signal is available.
Currently, RL is generally employed to further enhance the model’s accuracy in a specific domain by setting reward signals in a reward model to guide the model’s training [18], thereby achieving further performance improvement.
6 Performance in Speech Tasks
With the evolution of Speech LLMs, the integration of large language models (LLMs) into speech-based systems has yielded substantial advancements across various dimensions. This section investigates the current performance of Speech LLMs in several pivotal tasks in speech understanding, evaluating how LLMs enhance the capabilities of speech models compared to traditional methods. We will primarily focus on two classic speech-related tasks: Automatic Speech Recognition(ASR), and Speech Translation(ST), each epitomizing a crucial aspect of speech applications. Meanwhile, we will further explore the multi-tasking and cross-tasking capabilities exhibited by Speech LLMs across various tasks in the speech understanding field.
6.1 Automatic Speech Recognition
In this survey, we assess the performance of various models via the Word Error Rate (WER) tested on the clean and other test set of LibriSpeech dataset [46], which is an essential benchmark for evaluating ASR systems.
Stage Models
test-clean
test-other
Pre-LLM Stage Deep Speech 2 3.5 10.6
Early LLM Stage Conformer 1.9 3.9
HuBERT Base 3.4 8.1
HuBERT X-Large 1.9 3.5
Whisper Large-V2 2.5 4.9
Whisper Large-V3 1.8 3.6
Current Speech LLM Stage Qwen-audio 2.0 4.2
SALMONN 2.1 4.9
Seed-ASR 1.6 2.8
Table 1: Model Performance on ASR task (WER%)
Fig. 5: Model Performance on ASR task (WER%)
Traditional models based on LSTM or GRU architectures have consistently demonstrated strong performance across various benchmark datasets. For instance, the LSTM-based model Deep Speech 2 achieved a Word Error Rate (WER) of approximately 3.51% on the LibriSpeech clean test set and 10.62% on the other test set [22]. The introduction of Large Language Models (LLMs) has significantly improved these results. Notably, the Conformer model reduced the WER to 1.9% and 3.9% on the clean and other test sets, respectively [8]. Similarly, the HuBERT Base model, trained solely with 960 hours of LibriSpeech audio, achieved WERs of 3.4% and 8.1%. Utilizing an extended dataset of 60,000 hours from Libri-light, the HuBERT X-Large model further reduced the WERs to 1.9% on the clean test set and 3.5% on the other test set. As of 2023, further advancements were seen with Whisper Large-V3, achieving a new low WER of 1.8% on the clean test set. Recent Speech LLM models like Qwen-audio [17] and SALMONN [16] have also shown impressive results, approaching state-of-the-art performance, with WERs around 2% on the clean test set and below 5% on the other test set. Currently, the most advanced ASR system on the LibriSpeech dataset is Seed-ASR, which achieves a WER of 1.6% on the clean test set and 2.8% on the other test set [18].
6.2 Speech Translation
Fig. 6: Model Performance on speech translation and machine translation task (BLEU Score)
In the realm of natural language processing, translation tasks, encompassing both speech translation (ST) and machine translation (MT), are pivotal for enhancing global communication. Just like other NLP fields, these translation tasks have experienced significant progress due to the recent advancements in large language models (LLMs) [47]. In this survey, we employ the BLEU (Bilingual Evaluation Understudy) score to evaluate the translation performance of each model on the two-way translation between Chinese and English. For speech translation, we utilize the FLEURS test sets [48], and for machine translation from Chinese to English, we use the FLORES test sets [49, 50]. Additionally, the WMT’20 test sets [51] are used for English-to-Chinese machine translation.
There are two primary objectives in the translation domain: enhancing translation quality and expanding language coverage. In machine translation, the NLLB model, which is fine-tuned on LLMs aiming to be a universal translation system, initially broadened its linguistic scope to include more than 200 languages [50]. Furthermore, BigTranslate, leveraging instruction tuning based on LLaMA, has expanded the translation applicability of LLMs to a broader array of languages, achieving results comparable to those of ChatGPT and Google Translate [27]. The advent of LLMs has also propelled advancements in speech translation. Using a Cascaded ASR + MT approach, Whisper has shown notable efficacy in speech translation by utilizing extensive web-scale data [23]. Subsequently, AudioPaLM2 [14] enhanced speech translation quality by integrating text and speech-based models, thereby achieving impressive results as an end-to-end speech translation model. Previously achieving state-of-the-art in both ST and MT tasks on various public datasets, SeamlessM4T [29], a foundational multilingual and multitask model, is now surpassed by GenTranslate. This model utilizes the information stored in N-best candidates instead of relying solely on beam search decoding’s top-1 selection, a common practice among other models [28]. Please refer to  Figure 6 for a detailed performance comparison of these models.
6.3 Multi-tasking and Cross-tasking Abilities
Besides the above-mentioned two speech tasks, some Speech LLMs also demonstrate remarkable results on a wide range of speech tasks, showcasing their good multi-tasking and cross-tasking capabilities. Among these, Pengi is notable for its versatility, as it supports both open-ended tasks like Audio Captioning and Audio Question Answering (AQA), and close-ended tasks such as Sound Event and Scene Classification, without requiring any task-specific fine-tuning or extensions. Pengi’s approach involves framing all audio tasks as text-generation tasks, leveraging a unified architecture that effectively blends audio and text inputs into text outputs, achieving state-of-the-art performance on several benchmarks, including audio question answering and sound event classification [15].
Following Pengi, SALMONN further broadens these capabilities by integrating both speech and general audio inputs, such as music, into a single multimodal model. This integration enables SALMONN to perform well in emergent tasks that were not explicitly trained for, such as Audio-based Storytelling and Speech Audio Co-reasoning, illustrating its comprehensive cross-modal understanding. Its novel architecture allows it to handle complex tasks across different modalities, showcasing the potential for multimodal AI systems that can ”hear” and comprehend a wide range of auditory information [16].
Lastly, Qwen-Audio significantly advances the field of audio interaction by scaling up audio-language pre-training across more than 30 tasks involving various audio types. This model excels in areas like Acoustic Scene Classification, Speech Emotion Recognition, and Automatic Audio Captioning. By employing a multi-task training framework that minimizes interference from the diversity of tasks and audio types, Qwen-Audio achieves groundbreaking results in these areas, indicating a significant enhancement in audio understanding capabilities [17].
7 Challenges
Despite significant progress in applying large language models (LLMs) to the field of speech recognition, especially in fixed tasks such as Automatic Speech Recognition (ASR), or leveraging the powerful capabilities of LLMs to accomplish some multitasking, numerous challenges and issues remain. We validated and tested some of the most cutting-edge current works, which helped to corroborate the problems we identified.
7.1 LLM’s Dormancy
Regarding the method of using a projector to project audio feature embeddings into the text token space and combining them with text token embeddings of prompts, although it shows high accuracy and considerable generalization ability in ASR tasks, the LLM does not respond positively to prompts that were not seen during training. We refer to this phenomenon as LLM dormancy. To further illustrate the issue, we use SLAM-ASR and SALMONN as replication examples:
The core idea of SLAM-ASR is projecting audio feature embeddings to text token embeddings through a projector composed of a linear neural network. Here, we use a recipe consisting of wavlm-large, a projector, and vicuna-7b, with whisper-large and vicuna-7b kept frozen during training while only the projector is trainable. We trained for three epochs on the librispeech-960h dataset and achieved results on the librispeech test_clean and test_other datasets close to the expectations outlined in the original paper.
The core idea of SALMONN is to use a window-level Q-former to integrate audio information extracted by BERT and Whisper and map it into the text token space. During training, we updated the parameters of the LLM and encoder using LoRA and the window-level Q-former. We achieved results on librispeech test_clean, test_other, and Gigaspeech [52] that matched the expectations of the original paper.
We conducted a series of prompt tests on the two models mentioned above, with the results as shown in Table 2. From the table, we can see that SLAM-ASR almost does not respond to prompts, while SALMONN selectively responds to prompts, which precisely confirms our view that LLMs cannot maintain full model performance in the current framework.
Input: Can you divorce life from art?(A man speaker with chinese accent)
Output
Ability Test Prompt SLAM-LLM SALMONN
ASR transcribe speech to english text Can you divorce life from art? can you divorce life from art?
Speech Translate transcribe the speech to text and translate it to chinese. {CJK}UTF8gbsn你能从艺术中离开吗?
Speech Translate transcribe the speech to text and translate it to Japanese. {CJK}UTF8minアートから生活を破棄することができますか?
Speaker Verify which scene will this talk occur? The speaker is male and has an American accent. He speaks with confidence and clarity.
Speaker Verify Is the speaker a female? Yes, the speaker is a female.
Dialogue Ability Tell me the reason. I’m sorry, I’m not sure what you’re asking. Could you please provide more context or clarify your question?
Dialogue Ability
Based on the speech, can you help me determine
the answer?
Yes, I can divorce life from art.
Dialogue Ability can you answer the speech and give reasons? Can you divorce life from art
Context Inference
Actually this is a speech a man asked me. I don’t
have answer. And someone tells me to say yes.
But I don’t know why. Can you explain it?
Can you divorce life from art?
Yes, I can explain it. The man asked you if you could
divorce life from art. This means that you can separate
the concept of life from the artistic representation of it.
For example, a painting of a landscape could be seen
as a representation of life, but not necessarily a reflection
of it. So, the answer to the man’s question is yes, you can
divorce life from art.
Table 2: Prompt Tests conducted on Slam-ASR and SALMONN; Text highlighted in red indicates errors or incorrect outputs, while text in black represents reasonable or correct responses.
From the perspective of training strategies, SLAM-ASR did not update parameters for the encoder or LLM, thus making it more intuitive that the problem likely arises from the intermediate modality fusion step. To validate our approach, we attempted to reproduce the training process of SLAM-ASR. We then compared the average absolute value of the projected audio feature vectors with the average absolute value of the token embeddings obtained from encoding the prompt.
Fig. 7: Mean absolute value of embeddings
Note: As the tensor shape of the input embedding is (Batch size, Sequence id, Feature dimension), we set batch size to 1 and calculate the average of the absolute value of all the elements corresponding to the sequence id.
The figure here is a randomly chosen audio input for clear demonstration.
From Figure 7, we observe that the audio embeddings are significantly larger compared to the text embeddings.
Referring back to the procedure in the transformer module in a typical large language model, the text is first converted into tokens and embeddings before calculating the attention mechanism, which includes positional embeddings. However, the audio feature vectors are projected and then directly concatenated with the text token embeddings in Speech LLM. Clearly, there is no attempt to capture the positional information of the two modalities here.
A second potential issue arises during the calculation of attention scores. Recalling the standard steps involved in calculating attention scores in the attention mechanism, Given the audio-text modality embedding vector $E$, and since the LLM is kept frozen, the weight matrices $W^{Q}$, $W^{K}$, and $W^{V}$ are fixed. We transform $E$ into query, key, and value vectors using linear transformations, calculated as follows:
$Q$ $=EW^{Q}$ (3)
$K$ $=EW^{K}$ (4)
$V$ $=EW^{V}$ (5)
Next, we compute the attention scores:
Attention $\text{Attention Scores}=QK^{T}$ (6)
where $K^{T}$ is the transpose of the key vector.
To stabilize the training and inference process, we optionally scale the dot-product attention scores:
Scaled $\text{Scaled Scores}=\frac{QK^{T}}{\sqrt{d_{k}}}$ (7)
where $d_{k}$ is the dimensionality of the key vectors.
The scaled attention scores are then transformed into probability distributions using the softmax function:
[…]
Finally, we use the attention weights to perform a weighted sum of the value vectors to obtain the final attention output:
Attention Output=Attention $\text{Attention Output}=\text{Attention Weights}\cdot V$
which in detail can be expressed as:
Attention $\text{Attention Output}=\left(\text{Softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)\right)\cdot V$ (9)
Here we can observe that since the large language model is frozen, only $E$ changes based on the input. The embedding $E$ is obtained by concatenating the audio feature embedding $A$ and the text feature embedding $T$:
$E=[A\;T]$
We can also decompose $Q$, $V$, and $K$ into vectors corresponding to $A$ and $T$, for example,
$Q=[Q^{A}\;Q^{T}]$
$V=[V^{A}\;V^{T}]$
$K=[K^{A}\;K^{T}]$
Since $A\gg T$, according to equations (1), (2), and (3), due to the shared weight matrices, $Q^{A}\gg Q^{T}$ and similarly for other vectors. It is clear that the final attention output from the speech part is much larger than that from the text part. As a result, the model is likely to neglect the text portion, leading to the large language model falling into dormancy.
This can also be explained more broadly and intuitively using Bayes’ theorem [16]. Let $Y$ denote the output, $X$ denote the audio feature, and $I$ denote the instruction.
$P(Y|X,I)=\frac{P(Y|X)\cdot P(I|Y,X)}{P(I|X)}$ (10)
If we provide only one or a few corresponding outputs $Y$ for a given audio input, then:
$\frac{P(Y|X,I)}{P(Y|X)}$ (11)
is not negligible. In other words, the influence of $I$ on the result will not be significant.
Models like SALMONN and Qwen-audio, which are Speech LLMs, perform poorly in tasks involving untrained speech emotion inference and environmental understanding [53], and may even provide irrelevant answers. This further supports the observation that we did not fully achieve modality alignment under the current framework.
7.2 High cost
The high usage costs can be divided into two main aspects: time cost and memory cost. Due to the complexity of the architecture and the large number of parameters in large language models, there is a significant delay during inference. For instance, models like Whisper that use the Transformer architecture are noticeably slower in inference compared to traditional models. Additionally, the large number of parameters imposes high demands on GPU memory, both during inference and training.
8 Future Exploration
According to the challenges described earlier, the primary issue to address is the alignment of text and speech modalities. Currently, expanding the token space of large language models to include audio tokens can achieve a thorough integration of modalities, enabling the transition from large language models to multimodal models [41]. However, this approach faces issues such as data imbalance and high training costs. Additionally, simply using connectors for fine-tuning to enable the model to understand mappings between continuous feature embeddings has not yielded the expected results. The core problem is that current model architectures and training paradigms cannot guarantee an increase in task accuracy while maintaining the performance of large language models. Researchers are beginning to explore alternative alignment methods, such as constraining or normalizing continuous features and converting them into discrete representations before mapping them to text tokens in large language models [37].
In terms of training strategies, innovation is also essential. Current training methods largely focus on pretraining and supervised fine-tuning, with reinforcement learning (RL) not yet being widely applied. However, when large language models are integrated into the model architecture, these training methods may not fully meet the requirements of training large models. Considering the diversity of inputs and outputs in large language models, exploring the incorporation of reinforcement learning strategies such as Proximal Policy Optimization (PPO) [54] can be beneficial. Introducing perturbations in supervised learning can help maintain the model’s multi-task performance. Additionally, to address hallucinations in large language models, techniques such as Reinforcement Learning with Human Feedback (RLHF) can be employed to apply ”soft” constraints [55], alleviating issues such as repetitive outputs and other hallucination phenomena.
Currently, large models are widely deployed in only a few specific tasks. Given the powerful contextual capabilities of large language models, researchers are beginning to explore their integration into more complex systems. For instance, there is a growing interest in incorporating large language models into dialogue systems [56], using them as the ”brain” to coordinate other components within the system. Researchers are also experimenting with integrating large language models into conference and translation systems [57]. The exceptional contextual understanding and reasoning capabilities of large language models can handle multi-state judgments and selections. More precisely, their outstanding performance has not yet been fully demonstrated across various domains. Considering latency and high training costs, it is more practical to utilize them as system coordinators.
Meanwhile, the potential of Speech LLMs in challenging areas such as long speech recognition and keyword detection is worth noting. Given the strong contextual capabilities of large language models, we might improve inference quality by segmenting or refining long speech text content and incorporating keywords and contextual information into the prompt for more comprehensive and high-quality reasoning.
9 Conclusion
This work presents a comprehensive exploration of Speech Large Language Models (Speech LLMs), focusing on their architecture, training strategies, and advancements in the field of Spoken Language Understanding (SLU). We examined the evolution from traditional methods to novel architectures that integrate Audio Feature Extraction, Multimodal Information Fusion, and LLM Inference, facilitating richer audio processing and deeper multimodal understanding. We summarized key methods for multimodal information fusion, including several approaches in audio-to-text modality conversion and combining audio and text feature spaces. Additionally, we discussed training strategies such as self-supervised learning and reinforcement learning that enable enhanced performance across various speech tasks. The survey highlights key improvements in Rich Audio Transcription and the potential for Cross-task Integration across SLU tasks. However, challenges such as the Dormancy of LLMs under certain conditions persist, underscoring the need for further innovation in training strategies and system design. By addressing these challenges and exploring future solutions, this work provides valuable insights into advancing Speech LLMs and their applications in multimodal contexts.

### v4

@@
 Abstract
-Large Language Models (LLMs) exhibit strong contextual understanding and remarkable multi-task performance.
-Therefore, researchers have been seeking to integrate LLMs in the broad sense of Spoken Language Understanding (SLU) field.
-Different from the traditional method of cascading LLMs to process text generated by Automatic Speech Recognition(ASR), new efforts have focused on designing architectures centered around Audio Feature Extraction - Multimodal Information Fusion - LLM Inference (Speech LLMs).
-This approach enables richer audio feature extraction while simultaneously facilitating end-to-end fusion of audio and text modalities, thereby achieving deeper understanding and reasoning from audio data.
-This paper elucidates the development of Speech LLMs, offering an in-depth analysis of system architectures and training strategies.
-Through extensive research and a series of targeted experiments, the paper assesses Speech LLMs’ advancements in Rich Audio Transcription and its potential for Cross-task Integration within the SLU field.
-Additionally, it indicates key challenges uncovered through experimentation, such as the Dormancy of LLMs under certain conditions.
-The paper further delves into the training strategies for Speech LLMs, proposing potential solutions based on these findings, and offering valuable insights and references for future research in this domain, as well as LLM applications in multimodal contexts.
+Large Language Models (LLM) exhibit strong contextual understanding and remarkable multitask performance.
+Therefore, researchers have been actively exploring the integration of large language models (LLM) into the domain of speech understanding, with a primary focus on a broad range of speech-to-text tasks.
+These include automatic speech recognition (ASR), speech-to-text translation (ST), speech emotion recognition (SER), and others.
+We refer to such models as Speech LLMs, which are typically built on a unified architecture following the pipeline of Audio Feature Extraction $\rightarrow$ Multimodal Information Fusion $\rightarrow$ LLM Inference.
+This approach enables richer audio feature extraction while facilitating end-to-end fusion of audio and text modalities, thereby achieving deeper understanding and reasoning from audio data.
+This paper elucidates the development of Speech LLMs, offering an in-depth analysis of system architectures.
+Through extensive research and a series of targeted experiments, the paper assesses the advancements in Speech LLMs and its potential for cross-task integration within the speech understanding field.
+Furthermore, it indicates key challenges discovered through experimentation, such as the dormancy of LLM under certain conditions.
+The paper further delves into the training strategies for Speech LLMs, proposing potential solutions based on these findings, and offering valuable insights and references for future research.
 1 Introduction
 The advancements in large language models within natural language processing have been particularly significant, especially in contextual understanding and multimodal technology.
-Models like GPT-4 [1], trained on extensive corpora, are now capable of processing complex contexts with exceptional accuracy.
-These models, with their vast parameters, excel in parsing contextually appropriate sentences and maintaining coherence over multiple conversational turns, which is crucial for tasks like dialogue systems, automatic summarization, and machine translation.
-By leveraging vast contextual information, these models manage long-text processing with unprecedented proficiency in contextual reasoning.
-In addition to the success in text-based tasks, large language models have also achieved remarkable success in multimodal tasks [2], particularly in integrating and processing inputs from diverse modalities like vision, speech, and text.
+Models like GPT-4 [1], trained in extensive corpora, are now capable of processing complex contexts with exceptional accuracy.
+In addition to success in text-based tasks, large language models have also achieved remarkable success in multimodal tasks [2], particularly in integrating and processing input from various modalities such as vision, speech, and text.
 Their ability to semantically link these different types of data has led to significant breakthroughs in areas such as visual question answering and image generation, as exemplified by models like CLIP [3] and DALL-E [4].
-This capability not only enhances performance in visual tasks by effectively fusing information from different modalities, but also demonstrates the models’ potential to expand into and improve upon other modalities.
-The way these models manage complex multimodal information processing highlights their adaptability and lays the groundwork for further innovations in multimodal understanding and integration.
-The typical Spoken Language Understanding (SLU) involves interpreting spoken language to extract meaning, intent, and relevant information beyond simple transcription [5, 6].
-In order to more comprehensively and coherently demonstrate the performance of Speech LLMs, in this article, we also include basic speech-to-text tasks, such as automatic speech recognition (ASR), within its scope.
-Related works have achieved remarkable performance across various domains, including voice-activated assistants, automated transcription, and real-time speech translation.
-Modern systems are adept at handling diverse accents and languages, improving efficiency and accuracy in workflows like medical transcription and customer service automation.
-However, challenges remain, particularly in long-form speech recognition and hotword/keyword recognition.
-Long-form recognition struggles with maintaining context over extended periods and handling speaker variability, often resulting in accuracy degradation and latency issues in real-time applications.
-Hotword detection, critical for wake word recognition, faces difficulties in noisy environments and balancing sensitivity and specificity, especially when hotwords are contextually similar to other phrases.
-The integration of hotword recognition into continuous speech processing without disrupting flow or increasing computational overhead remains an ongoing challenge.
-In this paper, our contributions can be summarized in the following three points:
+In the field of speech processing, large language models have gradually begun to be widely applied [5, 6].
+Instead of using Automatic Speech Recognition (ASR) technology to recognize cascaded large language models  [7, 8], the Audio Feature Extraction - Multimodal Information Fusion - LLM Inference has become the most widely adopted approach.
+Unlike other survey papers that might also focus on generative tasks in speech  [9], our work does not center on speech generation, but rather focuses primarily on their development and application in speech understanding tasks, including traditional tasks such as acoustic speech recognition and speech-to-text translation, as well as tasks such as speaker recognition, speech emotion recognition, and others.
+Speech LLMs have not only achieved significant success in the tasks mentioned above  [10, 11, 12], but also demonstrated impressive capabilities and potential in challenging tasks such as long-form speech recognition, hotword and keyword recognition, etc.
+[13, 14]
+In this paper, our contributions can be summarized in the following three points.
   1. 1.
-    The first in-depth and comprehensive survey analyzing Speech LLMs in the SLU domain.
-We will conduct an in-depth exploration of the development of Speech LLMs, explaining the paradigm of model architecture, and perform a comparative analysis with traditional speech models in SLU tasks.
+    The first in-depth and comprehensive survey that analyzes Speech LLMs.
+We will conduct an in-depth exploration of the development of Speech LLMs, explaining the paradigm of model architecture, and perform a comparative analysis with traditional speech models in speech understanding tasks.
   2. 2.
-    The first comparative analysis of training methods for aligning speech and text modalities in Speech LLMs architecture, with an emphasis on the potential development of Reinforcement Learning strategies such as DPO and PPO.
+    The first comparative analysis of training methods for aligning speech and text modalities in Speech LLMs architecture, as well as the first to provide an in-depth classification and summary of the training datasets for Speech LLMs.
   3. 3.
@@
 2 Recent Advances in Speech LLMs
+2.1 Advancements of Speech LLMs in Speech Understanding
+With the increasing integration of large language models into the speech understanding, the Speech LLMs paradigm has shown substantial progress across a range of speech understanding tasks.
+Rather than replacing existing solutions, Speech LLMs enhance them by combining rich audio representations with strong language reasoning capabilities.
+In traditional tasks such as automatic speech recognition (ASR) and speech-to-text translation(ST), Speech LLMs consistently achieve higher accuracy than conventional models [15, 16].
+The use of continuous representations enables these models to retain more fine-grained acoustic and contextual information, which is particularly beneficial in high-precision applications.
+Furthermore, Speech LLMs have demonstrated notable improvements in multilingual and low-resource scenarios, expanding the accessibility and robustness of speech technologies [17, 18].
+Beyond accuracy, another key advantage of Speech LLMs lies in their multitasking capability.
+A single model can seamlessly perform various speech understanding tasks, including ASR, translation, speaker identification, emotion recognition, voice activity recogintion and even keyword spotting  [19, 20, 21, 12].
+This unified modeling approach reduces system complexity and facilitates the development of more versatile speech systems.
+Moreover, Speech LLMs might address long-standing challenges in the field.
+In long-form speech recognition, they help maintain contextual consistency over extended durations, significantly reducing error rates in processing lengthy audio.
+In keyword and hotword recognition, they offer higher precision and better robustness in noisy or dynamic environments by leveraging global context and adaptive modeling  [14, 22, 23].
+Together, these advancements position Speech LLMs as a powerful foundation for building next-generation speech systems that are not only accurate but also adaptable and capable of handling diverse real-world tasks7.
 2.2 Evolution of Speech LLMs Architectures
-Considering the exceptional performance of large language models (LLMs) in contextual understanding, reasoning, and multimodal information processing, and recognizing the unresolved challenges in the field of SLU, such as long-form speech understanding and hotword recognition, researchers have begun to explore the integration of LLMs into SLU.
+We broadly categorize the development of Speech LLMs into two phases.
+The first phase corresponds to the precursors of Speech LLMs, where the transformer architecture, having demonstrated significant success in natural language processing (NLP), was subsequently applied to traditional speech recognition and related models.
+The second phase involves integrating large language models (LLM) as a unified module following the speech feature extractor, establishing them as a critical component of speech processing.
 2.2.1 Transformer Integration into Traditional Speech Models
-One of the key milestones in the integration of LLMs into SLU was the introduction of the Transformer architecture.
+Early explorations of integrating Transformers into speech models retained the traditional speech modeling paradigm, incorporating the Transformer architecture as a component rather than fundamentally altering the overall model structure.
 Dong et al.
-(2018) first applied the Transformer to speech recognition, proposing an end-to-end speech recognition system based on the Transformer [7].
-This attempt laid the foundation for subsequent research.
+(2018) applied the Transformer to speech recognition, proposing an end-to-end speech recognition system based on the Transformer [24].
 Following this, in 2020, Gulati et al.
@@
 In 2021, Hsu et al.
-proposed HuBERT (Hidden-Unit BERT, which represented a major breakthrough in utilizing LLMs to process audio features [9].
+proposed HuBERT (Hidden-Unit BERT) [26].
 HuBERT employs self-supervised learning on a large corpus of unlabeled speech data to learn effective hidden unit representations.
+The encoder part of the transformer processes the audio features and generates contextualized representations, which are then used for the prediction task.
 In 2022, Radford et al.
-introduced the Whisper model, which further integrated multilingual and multitask capabilities into a single model, demonstrating excellent transferability across different languages and tasks [3].
-Researchers, unsatisfied with simply using speech as input and text as output, began exploring end-to-end speech-text multimodal inputs and outputs.
-In 2022, Junyi Ao et al.
-introduced SpeechT5, pioneering the investigation of a unified encoder-decoder framework for various spoken language processing tasks [10].
+introduced the Whisper model, which further integrated multilingual and multitask capabilities into a single model.
+Whisper also leverages the power of the transformer architecture to handle the complex mappings between speech and text.
+The self-attention and cross-attention mechanisms of the transformer allow Whisper to attend to different parts of the speech signal when generating the text output, improving its accuracy and robustness [3].
 2.2.2 Direct Audio Processing with Large Language Models
 Fig.
-1: The evolution of Speech LLMs from the Pre-LLM Stage with separate acoustic and language models, to the Early LLM Stage with transformer integration, and the Current Speech LLM Stage, featuring direct audio processing by large language models.
+1: The evolution of Speech LLMs from the Pre-LLM Stage with separate acoustic and language models, to the Early LLM Stage with transformer integration, and the Current Speech LLMs Stage, featuring direct audio processing by large language models.
 The current stage includes two branches: Discrete Sequence Modeling (e.g., VALL-E, SpeechGPT) and Continuous Sequence Modeling (e.g., Pengi, SALMONN).
-Besides directly incorporating the architecture of LLMs like transformer into speech recognition tasks, another significant integration approach involves using LLM as a whole to process audio features extracted by traditional speech recognition tools.
-This approach leverages the powerful contextual understanding and reasoning abilities of LLMs to improve the accuracy and robustness of speech recognition and deeper speech understanding, and it has gradually evolved into the main trend of using large language models in the field of speech recognition.In other words, this refers to the narrower definition of Speech LLMs .
-The earliest attempts to align speech and text modalities by passing extracted features from speech to downstream language models can be traced back to 2020.
-Qi Liu et al.
-were the first to explore this approach, separating the mapping of speech feature information from the speech feature extraction process into two independent modules and introducing the LSTM model into the process of mapping speech feature information to text [11].
-This approach differs from traditional CTC-based end-to-end systems by isolating the acoustic feature-to-text mapping as a standalone unit.
-This separation laid the groundwork for the subsequent integration of large language models into this area, making it possible to embed LLMs more effectively in speech recognition tasks.
-Currently, Speech LLMs can mainly be divided into two categories 4.
-The first is Discrete Sequence Modeling, represented by approaches that condense audio feature information into discrete tokens, which are then passed to the LLM for processing.
-In early 2023, Microsoft released Vall-E, a typical study utilizing LLMs to process audio features.
-Vall-E introduces the Transformer architecture, combining audio features with the capabilities of language models to achieve more natural text-to-speech generation [12].
-Following this, the emergence of SpeechGPT marked a deep integration of speech models and LLMs, capable not only of processing audio but also interacting through natural language, bringing a new interactive paradigm to the field of speech recognition [13].
-In mid-2023, Google launched AudioPaLM, which further expanded the capabilities of speech recognition into multimodal processing, enhancing performance in multimodal tasks [14].
-The second is Continuous Sequence Modeling, where audio feature information is projected into continuous input embedding vectors and then transmitted to the LLM for processing.
+Besides directly incorporating the architecture of LLM like the transformer into speech recognition tasks, another significant integration approach involves using LLM as a whole to process audio features extracted by the audio feature extractor.
+This approach leverages the powerful contextual understanding and reasoning abilities of LLM to improve the accuracy and robustness of speech recognition and achieve a deeper understanding of speech, and it has gradually evolved into the main trend of using large language models in the field of speech understanding.
+In other words, this refers to the narrower definition of Speech LLMs .
+Over time, researchers have converged on two primary strategies for integrating speech with large language models (Figure 4).
+The first, known as discrete sequence modeling, involves converting raw audio into a sequence of discrete units.
+These units, which capture essential features of the speech signal, allow the LLM to operate in a familiar token-based paradigm, bridging the gap between speech and text processing.
+Importantly, these discrete tokens are not only suitable for language modeling but also enable high-quality reconstruction of the original audio through neural vocoders.
+In early 2023, Microsoft released Vall-E, a zero-shot text-to-speech model exemplifying how EnCodec’s tokens are used in a Speech LLMs.
+By modeling acoustic tokens instead of waveform samples directly, VALL-E can generate high-quality speech that preserves the speaker’s identity, emotion, and acoustic environment from the prompt  [27].
+Following this, the emergence of SpeechGPT marked a deep integration of speech models and LLM, capable not only of processing audio but also of interacting through natural language.
+By training on paired unit sequences and transcripts, the LLM learns to treat discrete HuBERT tokens as another language it can understand and generate.
+This allows SpeechGPT inherently “speak” and “listen”: the LLM can take spoken instructions that are converted to units and generate either a text or a sequence of units as a response [28].
+In mid-2023, Google launched AudioPaLM, which further expanded the capabilities of speech recognition into multimodal processing by leveraging another major line of work in speech tokenization from the wav2vec family, improving performance in multimodal tasks [29, 30].
+The second paradigm is Continuous Sequence Modeling, where audio features are projected into continuous embedding vectors that are directly consumed by the LLM.
+Compared to the discrete formulation, this approach offers a more direct and arguably more natural way to model speech, as it preserves the continuous nature of audio signals throughout the processing pipeline.
 In 2023, Deshmukh et al.
-introduced Pengi, which projects speech modality information into the text space of large models without altering any parameters in LLM such as embedding vector spaces [15].
-Subsequently, continuous audio feature information began to be used to convey richer representations to the LLM [16, 17, 18], resulting in significant performance improvements across a range of tasks in the SLU domain.
-In 2024, Ma et al.
-further proposed SLAM-LLM, which involves the addition of a linear projector, allowing the task to be greatly completed by training only the projection layer [19].
-Once researchers achieved significant improvements in recognition accuracy compared to traditional speech recognition technologies, they began exploring and attempting to address the issues inherent in these conventional methods.
-In 2024, Fathullah et al.
-introduced the Conformer mechanism, achieving advancements in handling long speech sequences [20].
-In the same year, SpeechX achieved breakthroughs in multilingual, multitask speech recognition, enabling seamless switching between languages and supporting challenges like long-form speech understanding and hotword recognition [21].
-2.2 Advancements of Speech LLMs in Key Tasks and Challenges
-With the continued advancement and application of large language models (LLMs) in the field of speech processing, the Speech LLM paradigm has demonstrated significant success across various tasks, often surpassing the performance of traditional models in many areas.
-Our primary focus is on traditional key tasks within the domain of speech understanding, which include automatic speech recognition (ASR), speech translation, and speaker identification, among others.
-A more detailed comparison and analysis will be presented in  Section 6.
-At the same time, we also focus on the exploration and breakthroughs of Speech LLM in challenging tasks such as long-form speech recognition, as well as keyword and hotword detection.
-Of course, the powerful contextual reasoning capabilities of large language models provide an alternative solution for more complex system tasks.
-In this regard, we have also explored areas such as real-time multimodal interaction.
-This section investigates both the performance of Speech LLMs on these standard tasks and how they address long-standing challenges in the field.
-2.2.1 Improvement in Traditional Tasks in Spoken Language Understanding
-Traditional tasks in the field of speech understanding include automatic speech recognition, speaker identification, and speech translation.
-The Speech LLM paradigm has demonstrated impressive performance in all these areas.
-Here, we will firstly take ASR as an example to briefly examine and understand the progress Speech LLM has made in this field.
-Automatic Speech Recognition (ASR) is the technology that enables machines to convert spoken language into text.
-Modern ASR systems, enhanced by LLMs, aim to achieve higher accuracy, better noise resilience, and greater adaptability to diverse accents and dialects compared to their predecessors.
-This task is foundational for voice-controlled applications, interactive voice response systems, and automated transcription services.
-One of the key metrics for evaluating speech recognition systems is the Word Error Rate (WER).
-Traditional models, such as those based on LSTM or GRU, have shown competitive performance on various benchmark datasets [22].
-However, the introduction of LLMs has significantly improved these results [23, 24].
-Moreover, in multilingual speech recognition, LLMs have demonstrated superior performance across various languages.
-On the Multilingual LibriSpeech (MLS) dataset, Conformer-based models have achieved significant reductions in WER across multiple languages [25].
-Notably, in low-resource languages such as French and German, great WER reductions have been reported [23].
-Speech translation is the process of converting spoken language from one language into written or spoken text in another language.
-It typically involves two key steps [26]: automatic speech recognition (ASR), which transcribes spoken words into text, and machine translation (MT), which translates the transcribed text into the target language.
-This technology is used in real-time applications such as multilingual meetings, conferences, and live broadcasts, enabling seamless cross-lingual communication.
-With the successful application of large language models in the field of machine translation [27], the speech translation domain has also begun to gradually incorporate large language models, aiming to achieve similar significant results.
-Advances in large language models have significantly improved the accuracy and fluency of speech translation systems.
-In the field of speech translation, the integration of large language models has significantly enhanced performance.
-Not only has it improved the accuracy of speech translation tasks [28], but it has also broadened the range of supported languages [29], making these models an invaluable addition to the field.
-The incorporation of large language models brings a complementary advantage, enriching both precision and diversity in multilingual speech translation.
-Additionally, LLMs have excelled in multitask learning scenarios.
-For example, the Qwen-Audio model has shown impressive performance in tasks that combine speech-to-text with other modalities, such as sentiment analysis and speaker identification, reducing WER by 10% and improving sentiment recognition accuracy by 8% compared to single-task models [17].
-For a detailed comparison and analysis of Speech LLMs’ performance in traditional speech tasks, please refer to Section 6.
-2.2.2 Long-Form Speech Understanding
-Traditional speech recognition systems have struggled with long-form speech understanding due to context loss over extended periods.
-This issue is particularly pronounced in audio segments longer than one minute, where traditional models often exhibit a sharp increase in WER.
-In contrast, LLM-based systems have significantly mitigated this problem.
-The Whisper model, for instance, maintains contextual consistency across long-form audio, demonstrating an 18% reduction in WER on audio segments exceeding five minutes compared to traditional models.
-This enhancement is critical for applications such as the transcription of speeches and lectures.
-Additionally, models like UniAudio and Pengi have shown remarkable performance in maintaining low WERs across extended speech segments by integrating advanced contextual understanding [30, 31].
-These models have been particularly effective in domains requiring detailed, accurate transcriptions of prolonged speech events, where maintaining context over extended durations is crucial.
-2.2.3 Hotword Recognition
-Hotword recognition has been another challenging area for traditional speech recognition systems, especially in noisy environments.
-Traditional models often fail to accurately detect and differentiate hotwords within long audio streams.
-The GenTranslate model, introduced in late 2023, has made substantial improvements in this area by leveraging the contextual understanding capabilities of LLMs [28].
-In a test set containing multiple domain-specific hotwords, GenTranslate achieved a 22% improvement in hotword recognition accuracy compared to traditional models.
-Moreover, it maintained high robustness in noisy conditions, marking a significant advancement over prior approaches.
-Further examples of LLM-based models outperforming traditional systems include Mala-ASR and Whisper, which not only improves hotword recognition accuracy [32] but also adapts dynamically to new hotwords in real-time, a feature that traditional models lack.
-This capability is particularly valuable in dynamic environments like live broadcasts or interactive voice response (IVR) systems where new terms may frequently be introduced.
-2.2.4 Real-time Multimodal Interaction
-The integration of LLMs into speech recognition has expanded the scope of tasks beyond traditional speech-to-text, enabling real-time multimodal interaction.
-Models like VoxtLM and LauraGPT facilitate seamless integration of speech with visual and textual inputs [33, 34], providing coherent and accurate multimodal outputs.
-This capability is particularly valuable in applications such as live transcription and synchronized translation during presentations, where both speech and visual context need to be processed simultaneously.
-Moreover, LLM-based systems have introduced new functionalities, such as the generation of descriptive text, summaries, and even translations based on audio input.
-For instance, ViOLA can generate coherent summaries and cross-language translations with high fluency and accuracy, outperforming traditional models in both speed and quality of output.
-This represents a significant advancement in how speech recognition systems can interact with and interpret complex multimodal data streams [35].In addition, the integration of large language models with acoustics has demonstrated significant development potential across various specialized fields [36].
+introduced Pengi, which projects information about speech modality into the text space of large models without altering any parameters in LLM such as embedding vector spaces [31].
+By preserving more of the rich information inherent in the audio signal, continuous embeddings enable Speech LLMs to achieve state-of-the-art performance on high-precision tasks such as ASR and ST [10, 32].
+This paradigm’s ability to retain fine-grained acoustic details makes it particularly well-suited for applications where accuracy is critical.
+Moreover, these models also exhibit strong multitask capabilities, enabling them to generalize across various speech-related tasks [20, 19, 21].
+This approach might require only the insertion of a simple projection layer between the audio encoder and the LLM and a minimal fine-tuning cost, yet it can achieve state-of-the-art performance on certain tasks [33, 15].
+Researchers have not only pursued multitask capabilities and higher recognition accuracy beyond what traditional speech recognition technologies offer, but have also simultaneously sought to address long-standing challenges inherent in conventional speech understanding frameworks using this architecture [34, 14].
+Considering the focus on the field of speech understanding, the second approach is more widely adopted.
+Therefore, this paper will primarily elaborate on the second approach while providing a comparative analysis with the first approach.
 3 Model Paradigm of Speech LLMs
 As a model architecture that integrates both audio and text modalities, Speech LLMs exhibits significant differences in input-output processing and model structure compared to traditional speech models.
-Notably, due to the significant commonality and prevailing trends, we choose to focus on the narrower definition of Speech LLMs 1, which involves the integration of LLMs as a whole into SLU tasks.
-3.1 Speech LLM Input-Output Modes
-Fig. 2: Overview of Speech LLM architectures corresponding to three different input-output modalities: Speech-to-Text (S2T), Speech&Text-to-Text (ST2T), and Speech&Text-to-Speech&Text (ST2ST).
-Generally, the input and output modes of Speech LLMs can be categorized into three main types:
+Notably, due to the significant commonality and prevailing trends, we choose to focus on the narrower definition of Speech LLMs 1, which involves the integration of LLM as a whole into Speech Understanding tasks.
+3.1 Speech LLMs Input-Output Modes
+Fig. 2: Overview of Speech LLMs architectures corresponding to three different input-output modalities: Speech-to-Text (S2T), Speech&Text-to-Text (ST2T), and Speech&Text-to-Speech&Text (ST2ST).
+Speech LLMs can generally be categorized into three main types based on their input and output modes, with a particular focus on speech understanding tasks in this paper.:
   1. 1.
@@
     Speech&Text to Text (ST2T)
-    The ST2T mode, the most widely used in modern speech recognition, incorporates both audio and text inputs.
+    The ST2T mode, the most widely used in speech understanding field, incorporates both audio and text inputs.
 Typically, the text input serves as an instruction or task prompt.
 The model processes both modalities, fusing the audio and text information to produce a final text output.
-This approach enables multi-task capabilities and allows models to fully leverage the strengths of large language models, handling a wider variety of tasks than the S2T mode.
+This approach enables multitask capabilities and allows models to fully leverage the strengths of large language models, handling a wider variety of tasks than the S2T mode.
+This is also the primary mode of focus in this paper.
   3. 3.
     Speech&Text to Speech&Text (ST2ST)
-    In this more advanced mode, a vocoder is added to the output, enabling the model to generate both text and audio outputs.
-This expands the model’s capabilities beyond basic speech recognition, supporting tasks such as Text-to-Speech (TTS), speech translation, and voice conversion.
-The ST2ST mode offers greater versatility than the other two modes, making it applicable across a broader range of speech and language tasks.
-3.2 Model structure
-Fig. 3: Overview of key stages of the Speech LLM Workflow
-Up to now, various Speech LLM architectures have been developed, all of which are structured around three fundamental stages: Modality Feature Extraction, Modality Information Fusion, and LLM Inference.
-As illustrated in Figure 3, despite different input and output modalities, the major stages that a model went through is consistent.
-Speech LLMs process input from both speech and text modalities using separate extractors tailored to the characteristics of each modality.
-The raw audio input is processed by an audio encoder, which extracts relevant acoustic features, while the text input is handled by the encoder of the large language model to capture linguistic features.
-Once these features are extracted, they are fused through various techniques, which will be elaborated upon in the following sections.
-After feature fusion, the combined multimodal information is sent to the large language model for final decoding and output generation.
-In the following sections, we will examine three distinct model architectures, each corresponding to a different input-output modality.
-(a) Direct Projection
+    Currently, the ST2ST mode is widely applied in text-audio multimodal interaction scenarios through discrete sequence modeling.
+In simple terms, this mode requires the model to perform three key stages: speech/text understanding, text-based response generation, and speech/text output.
+Consequently, it demands a broader range of capabilities, such as Text-to-Speech (TTS) and Speech Translation (ST).
+Although ST2ST is not entirely within the scope of speech understanding, it still has significant relevance to this field.
+Therefore, this paper includes an analysis and comparison of this mode.
+4 Model structure
+Fig. 3: Overview of key stages of the Speech LLMs Workflow
+Various Speech LLMs architectures have been developed up to now, all of which are structured around three fundamental stages: Modality Feature Extraction, Modality Information Fusion, and LLM Inference.
+As illustrated in Figure 3, despite differences in input and output modalities, the major stages that each model undergoes remain consistent.
+All Speech LLMs begin by encoding the input information, regardless of the modality.
+In addition to S2T models, which take only audio as input, other models leverage pretrained audio and text encoders to obtain the representations of both modalities.
+The encoded audio and text embeddings are then fused to prepare them for input into the language model.
+This step transforms the outputs of the feature extraction process into a format compatible with the language model’s input requirements.
+At present, most Speech LLMs focus on generating text output; therefore, the LLM inference process typically revolves around a text decoder that receives the fused multimodal features as tokenized input to produce text.
+For models that optionally support audio output, such as ST2ST-type models, an LLM decoder is trained to generate audio tokens, which are subsequently converted into speech using a vocoder.
+4.1 Modality Feature Extraction
+As mentioned in the previous chapter, during the initial feature extraction stage, the audio is processed in two distinct ways.
+These approaches primarily differ in the output representation format of the audio features.
+Continuous Sequence Modeling, which is more commonly adopted in current Speech LLMs, encodes audio using a pre-trained and fine-tuned audio encoder, producing an audio embedding of shape [batch_size, sequence_length, feature_dimension].
+In contrast, Discrete Sequence Modeling discretizes the input audio signal into a sequence of discrete audio tokens—comparable to text tokens derived from raw text—which are not represented in embedding form.
+In this section, we summarize common models and techniques used to extract audio features using these two approaches.
+4.1.1 Continuous Sequence Modeling
+Continuous sequence modeling is generally more straightforward and typically encodes audio in the form of a time-frequency representation extracted using traditional signal processing-based feature extraction methods from the raw audio signal.
+Given an audio input $X_{a}$, a pre-trained audio encoder is applied to generate the audio features:
+$Z_{a}=g(X_{a})$ (1)
+where $g$ denotes the audio encoding process.
+The most commonly used audio encoders in mainstream Speech LLMs are Whisper [18] and Conformer [25].
+Whisper is a sequence-to-sequence Transformer model trained on a wide range of speech processing tasks, and its encoder is recognized as one of the most powerful for audio input processing.
+Conformer is another widely adopted architecture that combines convolutional layers with Transformers to effectively model both local and global dependencies in speech signals, particularly for automatic speech recognition tasks.
+Additional encoders used in mainstream Speech LLMs include WavLM [35], a predictive, self-supervised learning (SSL) pre-trained model.
+In this stage, aside from the audio encoder itself, subsampling modules are often applied to reduce the temporal resolution of the input features, thereby decreasing computational costs and enabling the model to focus on higher-level representations.
+Ultimately, the output audio feature embeddings have the shape [batch_size, sequence_length, feature_dimension], serving as the input for the next stage.
+4.1.2 Discrete Sequence Modeling
+Discrete sequence modeling converts raw audio into a sequence of discrete tokens that represent acoustic content and can typically be decoded back into high-quality audio.
+The core idea behind generating discrete tokens lies in vector quantization.
+Building on VQ-VAE [36], which introduced the concept of encoding continuous audio features into symbolic representations via a learned codebook, current mainstream methods for generating discrete tokens include self-supervised pretrained audio tokenizers such as HuBERT [26], and neural codec models such as EnCodec [37, 38].
+A general process involves first encoding raw audio into a sequence of latent representations, as in continuous sequence modeling: $Z_{a}=g(X_{a})$, where $Z_{a}=[\textbf{z}_{1},\textbf{z}_{2},\dots,\textbf{z}_{T}]$ with $\textbf{z}_{t}\in\mathbb{R}^{d}$.
+Next, a quantization function $q(\cdot)$ is applied to obtain the final discrete audio tokens:
+$u_{t}=q(\textbf{z}_{t})\ \text{with}\ \textbf{e}_{u_{t}}=\mathcal{Q}(\textbf{z}_{t})$ (2)
+Here, $\mathcal{Q}$ denotes the quantization function that maps continuous vectors to codebook vectors $\textbf{e}_{u_{t}}\in\mathbb{R}^{d}$, resulting in $u_{t}$, the discrete token index from the vocabulary set.
+The final output is a sequence of discrete audio token indices: $U=[u_{1},u_{2},\dots,u_{T}]$.
+Specifically, self-supervised pretrained audio tokenizers, such as HuBERT, learn contextualized audio representations via masked prediction tasks and discretize intermediate features using k-means clustering.
+The resulting cluster indices serve as phoneme-like discrete tokens.
+Neural codec models, such as EnCodec, adopt an encoder–quantizer–decoder architecture in which the audio is compressed into discrete tokens using residual vector quantization and later reconstructed [39].
+These models are trained end-to-end and produce high-fidelity audio that is well-suited for generative tasks.
+The resulting discrete audio tokens are processed alongside tokenized text, with the integration method detailed in the following section.
+4.2 Modality Information Fusion
+After completing the audio feature extraction stage, the most critical challenge in Speech LLMs lies in the alignment between the audio modality and the text modality.
+In this section, we discuss alignment methods in both continuous and discrete fashions.
+As introduced in the previous section, the output formats of these two methods differ, thereby requiring customized approaches to modality fusion in each case.
+4.2.1 Modality Fusion of Continuous Audio Embedding
+Before addressing how to effectively combine the two types of information, one important consideration is determining which specific part of the audio modality should be used.
+Researchers currently tend to use the output from the final layer of the encoder as the primary source of audio modality information.
+However, various alternative methods also exist.
+For example, some approaches utilize intermediate layer outputs to capture more granular features [40], while others apply attention mechanisms to emphasize relevant parts of the audio signal [35].
+Once the audio embedding is obtained, it is merged into the text embedding space using two main methods:
+  1. 1.
+    Pre-Decoder Alignment: This category refers to methods that align audio features with text embeddings before feeding them into the LLM decoder.
+A common approach is to project the audio feature information into the LLM’s text feature space through a connector [33, 20].
+Specifically, the tensor containing audio features typically has its own distinct feature dimension.
+The audio feature embedding $Z_{a}$ is projected into a vector with the same dimensionality as the text embedding using a projection function $w$: $H_{a}=w(Z_{a})$.
+These audio embeddings are then concatenated with the input text embedding $H_{t}$ to form a new embedding sequence that integrates both speech and text information: $H=[H_{a},H_{t}]$.
+This combined sequence is then fed into the LLM.
+Some researchers have also implicitly incorporated the projection step within the original encoder, achieving modality projection by adjusting the encoder’s parameters during training [19].
+Most models use a multi-layer perceptron (MLP) as the projection module, while some adopt more advanced techniques such as Q-Former, which introduces trainable query tokens that attend to audio features and output fixed-length embeddings aligned with the LLM input space [20].
+  2. 2.
+    In-Decoder Alignment (Cross Attention): In addition to merging modalities before inputting them into the language model, another method establishes a bidirectional interaction between audio and text features via cross-attention [41].
+In this approach, the model maintains separate audio and text streams, allowing each modality to attend to the other through cross-attention layers, thereby enabling deeper multimodal fusion within the transformer architecture.
+4.2.2 Modality Fusion of Discrete Audio Tokens
+For discrete audio tokens, they are typically treated in the same way as text tokens.
+The core challenge lies in how to process audio tokens using only a text-based vocabulary.
+There are essentially two ways to address this issue:
+  1. 1.
+    Token Mapping: The most intuitive approach is to map audio tokens to text tokens that the LLM can process.
+Tsunoo et al.
+[42] propose using CTC predictions to efficiently generate prompts—referred to as CTC prompts.
+Once text tokens are obtained, they are combined with tokenized input text to form a unified token sequence containing both audio and text information.
+This sequence is then fed into the LLM for processing.
+This approach not only preserves the semantic content of the audio features but also ensures consistency in the LLM’s processing pipeline.
+  2. 2.
+    Token Space Expansion: While projecting the speech modality into the text modality is straightforward, it does not achieve truly lossless modality fusion.
+Information loss and conflicts may occur during the modality conversion process (to be elaborated later).
+Therefore, researchers have proposed an alternative approach that modifies the input space of the large language model to natively incorporate the audio modality [30, 28, 43].
+Specifically, this method augments the token space by adding audio tokens to the existing text token vocabulary, forming a unified token space.
+These new audio tokens are synthesized from the audio features extracted in the previous stage, thereby preserving a greater portion of the original audio information (illustrated in Figure 4(c)).
+4.3 LLM Inference
+The LLM inference stage consists mainly of a text decoder, which is typically an autoregressive transformer decoder that generates text tokens conditioned on input audio and/or text.
+Architecturally, it often mirrors standard language models like GPT, but it is specially designed or adapted to handle multimodal inputs—such as projected audio embeddings or cross-attended audio features—alongside text tokens.
+To accommodate modality fusion, the decoder may incorporate special tokens (e.g., <aud>, <text>) or use prompt formatting to distinguish and contextualize different modalities during generation.
+Despite these adaptations, the decoder follows the same autoregressive objective: predicting the next text token given previous tokens and any additional input context.
+In ST2ST style models, the decoder may also be extended to support the generation of discrete audio tokens.
+One common approach, as seen in VALL-E, treats speech generation as a conditional codec language modeling task, where the model generates audio tokens conditioned on phoneme sequences and an acoustic prompt  [44, 27].
+Another approach used in SpeechGPT expands the LLM’s vocabulary to include both text and speech tokens, enabling the model to generate audio token sequences directly alongside or following text, depending on the instruction[28].
+In both cases, the generated audio tokens are passed through a neural codec decoder (e.g., EnCodec[37] or HiFi-GAN[45]) to reconstruct the final waveform.
+(a) Continuous Embedding Projection
 (b) Token Mapping
-(c) Combine audio and text space
+(c) Token Space Expansion
 Fig. 4: Illustration of three approaches to audio-text information fusion in Speech LLMs
-3.2.1 S2T
-S2T models (Speech-to-Text) take speech as the input and generate text as the output.
-These models typically incorporate an audio encoder to extract features from the speech signal, while a text encoder is not required, as text input is absent.
-Therefore, to optimize model size and efficiency, a decoder-only architecture is commonly employed to generate the text output.
-However, aligning the audio features with the input embedding space of the large language model (decoder-only architecture) requires a conversion module.
-This module ensures that the audio features are properly mapped, facilitating the generation of accurate text output.
-As discussed in the work by Tsunoo, et al.
-[37], the decoder-only architecture is particularly effective for streaming end-to-end speech recognition, where efficiency and real-time performance are critical.
-3.2.2 ST2S
-ST2S (Speech & Text to Speech) models process both speech and text inputs by employing distinct encoders for each modality during the modal information extraction phase.
-Specifically, an audio encoder extracts relevant features from the speech input, while a text encoder captures features from the textual input.
-In the subsequent modal information fusion phase, it is essential to align the speech-derived features with the textual representation space to ensure coherent integration.
-The unified multimodal features are then passed to the decoder of a large language model, which carries out the inference and decoding process to generate the final output.
-3.2.3 ST2ST
-Speech&Text-to-Speech&Text models (ST2ST) include an additional speech modality in the output, which requires a more complex decoding mechanism during the generation phase.
-After modal feature extraction, modal information fusion, and large language model inference, the text output can be generated by the decoder of the large language model.
-However, the speech output requires a vocoder to synthesize the final speech output [38, 13].
-4 Multimodal Information Fusion
-Having established a preliminary understanding of the overall architecture of Speech LLMs, we will now examine the model’s key components in greater detail.
-Beyond the well-known stages of audio feature extraction and large language model inference, the most critical issue in Speech LLM lies in the alignment between the audio modality and the text modality.
-Aligning these two modalities typically requires two steps, each addressing a specific problem: The first step, which we refer to as Audio Feature Post-Process, focuses on determining what specific audio modality information is needed.
-The second step, called Audio and Text Connection, addresses how to effectively combine these two types of information.
-4.1 Audio Feature Post-Process
-In this stage, researchers currently tend to directly use the final layer output of the encoder.
-The current mainstream approach is to extract the output of the final layer of the encoder as the primary source of audio modality information.
-However, there are other various alternative methods available.
-For instance, some approaches involve using intermediate layer outputs to capture more granular features [39].
-Others apply attention mechanisms to emphasize relevant parts of the audio signal [40].
-4.2 Audio and Text Connection
-Once the audio modality information is obtained, it must be integrated with the textual modality information to enable the LLM to perform the final inference.
-Currently, the methods used can be broadly classified into two categories: transforming the audio feature into the textual modality space, or merging the audio and textual modality spaces.
-4.2.1 Audio-to-Text Modality Conversion
-Converting speech modality information into the text modality space is one of the most commonly used approaches today.
-This method fully considers that large language models (LLMs) are primarily designed for text modalities, enabling direct projection and integration of the audio modality into the text modality space.
-This effectively minimizes modifications to the LLM while preserving its powerful capabilities.
-To achieve this, a connector or projector is needed to transform or project the extracted audio modality features.
-This neural network structure should minimize the loss of audio feature information while ensuring smooth transfer from the audio feature space to the text feature space.
-Currently, two common methods are employed (illustrated in  Figure 4(a) and Figure 4(b)):
+5 Training Strategies
+This section focuses on the training strategies of Speech LLMs mainly based on continuous embeddings, where speech signals are encoded into continuous feature vectors and directly consumed by a large language model.
+Training for these continuous-embedding Speech LLMs typically progresses through three major stages: Modality Alignment, Multitask Training, and Instruction and Preference Alignment.
+Each stage plays a distinct role in gradually aligning modalities, expanding task capabilities, and fine-tuning the model to follow human intent.
+5.1 Modality Alignment
+At the outset, Speech LLMs must learn a shared representation between spoken input and text.
+In this modality alignment phase, models typically leverage vast speech data to bridge the audio-text gap.
+A common strategy is to incorporate a pre-trained speech encoder and align its acoustic representations with a text-based LLM.
+For example, SALMONN and Qwen-Audio both begin with a pretrained Whisper encoder (originally trained for ASR) coupled to a LLM backbone [20, 12].
+The alignment is then refined by supervised tasks that force audio and text to correspond: notably, automatic speech recognition (ASR) and audio captioning (AAC) are usually used as fundamental alignment tasks  [46, 15].
+The model is trained to map an input audio sequence $a$ to a textual output $y$ (transcript or caption), effectively minimizing a cross-entropy loss
+$\mathcal{L}_{\text{align}}(\theta)=-\log P_{\theta}(y\mid a),$
+thereby tuning the interface between audio encoder and text decoder.
+In practice, only the bridging components (e.g.
+Q-Former modules or adapter layers) may be updated at this stage (the large language model and encoder often remain either frozen or are adapted with low-rank updates), ensuring the rich pre-trained knowledge in each modality is preserved while their features become synchronized.
+Self-supervised learning can further enhance alignment: some models (e.g.
+SEED-ASR) first pre-train the speech encoder on millions of hours of unlabeled audio to learn latent acoustic units, and then align these with text by transcribing massive paired data.
+Such pre-training may involve predicting masked audio segments or contrastively matching audio embeddings to pseudo-text units, building a modality-agnostic representation.
+The outcome of modality alignment is a foundation where spoken content is represented in the LLM’s latent space, capturing linguistic information from speech signals.
+This is evidenced by improved transcription and audio comprehension early in training—OSUM, for instance, reports that concurrently training an ASR task alongside any new audio task (an “ASR+X” strategy) accelerates alignment, as the model always learns to extract and decode the spoken words before handling additional subtasks.
+Overall, this phase imbues the model with a basic “speech understanding” ability, aligning phonetic and lexical content of audio with text tokens, which is crucial before tackling more complex or diverse tasks.
+5.2 Multitask Training
+With modalities aligned, the next phase develops broad audio-language capabilities via multitask training.
+Here, a single model is jointly trained on a wide variety of supervised tasks, spanning speech transcription, translation, and higher-level audio understanding.
+The training data consist of many triples $(a,p,y)$: an audio input $a$, a textual prompt or task descriptor $p$, and the expected textual output $y$.
+The model is optimized to predict $y$ from $(a,p)$ for all tasks in a unified fashion.
+Formally, if $\mathcal{T}$ is the set of task types, one can define a combined objective as the weighted sum of task-specific losses:
+$\mathcal{L}_{\text{multi}}(\theta)=\sum_{t\in\mathcal{T}}\lambda_{t}\;\mathbb{E}_{(a,p,y)\sim D_{t}}\big{[}-\log P_{\theta}(y\mid a,p)\big{]},$
+where $D_{t}$ is the training set for task $t$ and $\lambda_{t}$ balances the tasks.
+In practice, many Speech LLMs implement this by converting each task into an instruction-following format.
+For example, Qwen-Audio introduces special tokens or natural language prompts that indicate the task (<Transcribe>, <Translate>, “Describe the sound:”, etc.), so that the LLM can differentiate and perform over 30 diverse audio tasks.
+By co-training on multilingual ASR, speech-to-text translation, speaker identification, emotion classification, audio question-answering, music tagging, and more, the model acquires multi-faceted knowledge.
+This joint learning promotes generalization: shared acoustic-linguistic features are reused across tasks, and training signals from simpler tasks (e.g.
+phonetic recognition) assist more complex ones (e.g.
+spoken Q&A).
+In Qwen-Audio’s case, hierarchical task tags were used to encourage knowledge sharing while avoiding label conflicts, and the resulting model achieved state-of-the-art results on many benchmarks without task-specific fine-tuning.
+Academic models like OSUM followed a similar path with fewer resources: OSUM combined eight speech understanding tasks in training (from ASR to speaker attributes), always pairing each target task with an ASR objective.
+This ensured that the model never loses grounding in literal transcription, yielding more stable convergence and better overall performance.
+SALMONN likewise performed an instruction-based multitask fine-tuning (after initial alignment) on a wide range of speech, audio, and music understanding tasks.
+Notably, SALMONN’s multitask stage endowed it with numerous skills (recognizing speech, translating it, answering questions about audio, classifying emotions, etc.), effectively turning the LLM into a general audio assistant.
+A strength of this phase is the emergence of cross-task capabilities: since the model sees varied audio-text pairs, it can handle composite instructions and scenarios that were not explicitly in training.
+For instance, a multitask-trained model might learn to perform spoken language identification implicitly while doing translation, or combine speech recognition and reasoning to summarize an audio clip.
+By the end of multitask training, the Speech LLM has become a versatile audio-informed model with competencies across many domains, all encoded within one unified network.
+5.3 Instruction and Preference Alignment
+The final development phase focuses on aligning the model’s behavior with user instructions and preferences, ensuring that it follows natural language commands reliably and produces responses that are helpful and context-appropriate.
+This stage often begins with supervised instruction tuning.
+The model is fine-tuned on curated instruction-response datasets, where each example is a conversation or prompt $X$ (which may include an audio input and textual query) paired with a high-quality response $Y$.
+The fine-tuning objective is to minimize the negative log-likelihood of the reference response given the prompt:
+$\mathcal{L}_{\text{SFT}}(\theta)=-\mathbb{E}_{(X,Y)\sim D_{\text{instr}}}\left[\sum_{t=1}^{|Y|}\log P_{\theta}(y_{t}\mid y_{<t},X)\right].$
+This supervised fine-tuning (SFT) teaches the model to produce coherent, instruction-following answers in a conversational style.
+For example, Qwen2-Audio undergoes an SFT stage to become an interactive assistant (“Qwen2-Audio-Chat”), trained on multi-turn dialogues that involve both voice and text instructions.
+The effect is to instill the model with the format of polite, helpful responses, and the ability to handle follow-up questions or mixed-modality inputs (such as a user speaking a query in audio followed by a text instruction).
+SALMONN similarly leveraged instruction tuning on assorted audio-text tasks formatted as user prompts, which improved its adherence to queries; however, the authors found that this alone caused the model to sometimes over-fit to the tasks seen (producing, say, verbatim transcriptions even for novel instructions).
+To mitigate this, SALMONN introduced a brief activation tuning step: a few-shot fine-tuning on prompts requiring more creative, open-ended outputs (like audio-based storytelling), which helped restore the model’s generalization and “emergent” abilities beyond the narrow supervised tasks.
+After supervised alignment, some advanced Speech LLMs apply preference alignment using reinforcement learning  [12].
+This step optimizes the model’s responses according to human feedback or a learned reward function, refining qualities such as factual accuracy, relevance, or user satisfaction [47, 48].
+A prominent method is reinforcement learning from human feedback (RLHF) operationalized via Proximal Policy Optimization (PPO) [49].
+In RLHF, the model is treated as a policy $\pi_{\theta}(y|X)$ that generates a response $y$ to an input $X$, and a reward model $R(X,y)$ (trained from human preference rankings) scores the response.
+The goal is to maximize expected reward $E_{X}[E_{y\sim\pi_{\theta}}[R(X,y)]]$ while keeping the language model’s output distribution close to the pre-trained model to avoid degrading fluency.
+PPO achieves this by updating $\theta$ in the direction of the policy gradient $\nabla_{\theta}J\approx E_{X,y}[R(X,y)\,\nabla_{\theta}\log\pi_{\theta}(y|X)]$ with constraints on the KL-divergence from the original policy.
+In practice, the PPO update uses a clipped surrogate objective to ensure stable improvements.
+An alternative, lightweight approach is Direct Preference Optimization (DPO)  [50], which was employed in Qwen2-Audio’s final training.
+DPO sidesteps training a separate reward network by directly using preference pairs $(y^{+},y^{-})$ for a given prompt $X$.
+It adjusts the model to increase the probability of preferred response $y^{+}$ over the disfavored $y^{-}$.
+One can frame the DPO loss as:
+$\mathcal{L}_{\text{DPO}}(\theta)=-\mathbb{E}_{(X,y^{+},y^{-})}\left[\log\sigma\!\Big{(}\frac{1}{\beta}(\log\pi_{\theta}(y^{+}|X)-\log\pi_{\theta}(y^{-}|X))\Big{)}\right]\!,$
+where $\sigma$ is the sigmoid function and $\beta$ a temperature hyperparameter.
+Minimizing this loss pushes $\pi_{\theta}$ to assign higher probability to the superior response $y^{+}$, aligning the model with human preferences without the complexity of a reward model loop.
+By the end of this reinforcement or preference alignment phase, the Speech LLM is highly tuned to follow instructions in a human-preferred way: it will listen to a user’s voice request and respond with appropriate content, tone, and detail.
+Crucially, this stage often improves the model’s factuality and safety (by penalizing incorrect or undesirable continuations) – Qwen2-Audio’s authors note that after DPO fine-tuning, the model more faithfully adheres to user intent and exhibits fewer mistakes in complex audio-centric instructions.
+The strength of the instruction alignment phase lies in making the model not just capable, but also reliably usable: it transforms the multitask-trained foundation into a polite conversational agent that can handle open-ended voice interactions.
+However, it is worth noting that this phase has received limited attention in current Speech LLM research focused on speech understanding.
+Most existing works primarily emphasize modality alignment and task-specific performance, while instruction-following behavior, dialogue consistency, and safety alignment remain underexplored.
+As Speech LLMs continue to evolve toward general-purpose audio agents, further investigation into instruction and preference alignment is expected to play a crucial role in improving usability, controllability, and user interaction quality.
+All of these stages combined – modality alignment, broad multitask learning, and final alignment with instructions/preferences – produce state-of-the-art Speech LLMs that can understand speech in rich context and interact with users naturally, bridging the gap between spoken language and the powerful reasoning of text-based LLMs.
+(a) Combine audio and text space
+(b) Combine audio and text space
+Fig. 5: Two ways of generating SFT data for Speech LLMs training
+6 Datasests
+Speech Large Language Models (Speech LLMs) are increasingly benefiting from the expanding variety of datasets.
+In this section, we categorize and analyze these datasets from two perspectives: the source of the data and its structural composition.
+6.1 Datasets Based on Real-Recorded Audio
+6.1.1 Manually Annotated Datasets
+Traditional Single-Task Datasets
+A significant portion of the datasets used to train Speech LLMs consists of audio-text pairs.
+Numerous traditional single-task datasets are readily available, providing audio-text pairs for training.
+The most common manually annotated datasets that developed for traditional Speech tasks, include Automatic Speech Recognition (ASR), Speech Translation(ST), Speaker Diarization (SD), and other single-task-oriented datasets.
+These datasets are widely used across various branches of speech understanding due to their early development and large availability.
+Besides using these tasks to pre-train LLM, these datasets can also be directly used to train downstream tasks for finetuning the Large Language Models.
+The most commonly used datasets are as follows:
   1. 1.
-    Direct Projection: The audio feature information is directly projected into the LLM’s text feature space through a connector [19, 16].
-Specifically, the tensor containing audio features is projected into an embedding vector aligned with the text modality.
-These audio embeddings are then concatenated with the input text’s embedding vector to create a new embedding vector that integrates both speech and text information.
-This combined vector is then fed into the LLM.Of course, some researchers have also implicitly handled the projection step by internalizing it within the original encoder, achieving the modality projection effect by adjusting the encoder’s parameters during training [17].
+    Automatic Speech Recognition (ASR): LibriSpeech (English)[51], LibriLight (English)[52], GigaSpeech (English)[53], Aishell1 (Mandarin Chinese)[54], Aishell2 (Mandarin Chinese)[55], Voxpopuli (Multilingual)[56]
   2. 2.
-    Token Mapping: The audio feature information is mapped into text tokens that the LLM can process [37].
-Initially, the audio features are passed through a projector or converter to generate symbols or representations that correspond to text tokens.
-These representations are then combined with the tokenized input text, forming a token sequence that includes both audio and text information.
-This sequence is subsequently input into the LLM for processing.
-This approach not only preserves the integrity of the audio feature information but also ensures consistent processing within the LLM.
-4.2.2 Combining Audio and Text Feature Space
-Projecting the speech modality into the text modality is certainly straightforward, but it does not achieve lossless modality fusion in the true sense.
-Information loss and conflicts may occur during modality conversion (to be elaborated later).
-Therefore, researchers have proposed another modality alignment approach, which involves modifying the original input space of the large language model to integrate the audio modality [14, 13, 41].
-Specifically, this method augments the token space by adding audio tokens on top of the existing text tokens, creating a new token space.
-These new audio token are synthesized from the audio features extracted in the previous step, thereby largely preserving the original audio information (illustrated in Figure 4(c)).
-5 Training Strategies
-The training of current Speech LLMs mainly includes three approaches: pretraining, supervised fine-tuning (SFT), and reinforcement learning (RL).
-These approaches are used in different training stages and scenarios due to their varying principles and effects.
-5.1 Pretraining
-Pretraining involves training models on large-scale unlabeled data to enable them to learn broad, general knowledge.
-Typically, in Speech LLMs, speech encoders undergo pretraining on audio-text pairs to capture audio features.
-Common training strategies, including self-supervised learning (SSL) [18], are widely used.
-However, to better integrate speech encoders with LLMs, some researchers attempt to re-pretrain speech encoders [42], which allows for more direct acquisition of encoders with audio feature extraction capabilities.
-For Speech LLMs of Section 4.2.2, as mentioned previously, a thorough re-training of multimodal large models is necessary.
-This involves referencing the pretraining phase of large language models.
-By organizing mixed audio and text data, we adapt the pretraining methods of large language models to accommodate bimodal information.
-5.2 Supervised Fine-Tuning (SFT)
-For specific tasks in the speech domain, further fine-tuning is often required.
-Supervised fine-tuning is a common approach where high-quality labeled data from downstream task datasets is used to train the model, improving its performance on these tasks.
-In Speech LLMs, for models of structure as mentioned earlier in Section 4.2.1, supervised fine-tuning is an extremely common method.
-It is generally used to achieve alignment between the speech encoder and the LLM and to enhance performance on specific tasks.
-Common training methods include fine-tuning connectors, fine-tuning the encoder, and LLMs, such as using methods like LoRA.
-This mainly involves handling modality alignment and completing the model’s learning of text-token mapping.
-Similarly, models of structure described in  Section 4.2.2 also require supervised fine-tuning, which is essential for enhancing its performance on specific tasks.
-It is worth noting that supervised fine-tuning is also frequently used as a wake-up method for large language models.
-When modality alignment is completed, large language models might experience a certain degree of dormancy [16].
-Researchers use audio-based dialogue wake-up fine-tuning to reactivate large language models.
-5.3 Reinforcement Learning (RL)
-Reinforcement learning (RL) is a commonly used method in the training of large language models, especially in the field of large language model safety alignment [43].
-Unlike other training strategies, reinforcement learning ensures that the LLM optimizes in the desired direction while maintaining diversity in its outputs.
-In order to maintain and enhance cross-task integration capabilities of large language models under the Speech LLM paradigm, reinforcement learning is an effective training strategy that is worth further exploration [44, 45].
-However, it has not been widely used in Speech LLMs.
-Reinforcement learning is a framework where an agent interacts with an environment to maximize cumulative rewards over time.
-The agent learns by receiving feedback in the form of rewards, adjusting its actions to improve future rewards.
-Two commonly used RL algorithms in this context are Proximal Policy Optimization (PPO) and Direct Policy Optimization (DPO).
-Proximal Policy Optimization (PPO) is a policy gradient method that aims to optimize the objective function while limiting policy updates to avoid large deviations from the current policy.
-The objective function for PPO is defined as:
-$\mathcal{L}^{\text{PPO}}(\theta)=\hat{\mathbb{E}}_{t}\left[\min\left(r_{t}(\theta)\hat{A}_{t},\text{clip}(r_{t}(\theta),1-\epsilon,1+\epsilon)\hat{A}_{t}\right)\right]$ (1)
-where $r_{t}(\theta)=\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{\text{old}}}(a_{t}|s_{t})}$ is the probability ratio between the new policy and the old policy, $\hat{A}_{t}$ is the advantage function, and $\epsilon$ is a hyperparameter that controls the clipping range.
-The clipping mechanism ensures that policy updates remain within a trust region, preventing overly aggressive updates that could harm the model’s performance.
-Direct Policy Optimization (DPO), on the other hand, focuses on directly optimizing the policy by defining a reward function specific to the task. The objective in DPO can be expressed as:
-$\mathcal{L}^{\text{DPO}}(\theta)=\hat{\mathbb{E}}_{t}\left[R(s_{t},a_{t})\log\pi_{\theta}(a_{t}|s_{t})\right]$ (2)
-where $R(s_{t},a_{t})$ is the task-specific reward function, and $\pi_{\theta}(a_{t}|s_{t})$ is the policy parameterized by $\theta$.
-DPO directly maximizes the expected reward by focusing on optimizing the policy without clipping, often used in settings where a more direct reward signal is available.
-Currently, RL is generally employed to further enhance the model’s accuracy in a specific domain by setting reward signals in a reward model to guide the model’s training [18], thereby achieving further performance improvement.
+    Speech Translation: CovoST2 (21 source languages and 15 target languages)[57], Voxpopuli (14 languages)[56]
+  3. 3.
+    Automatic Audio Caption: Clotho (English)[58], AudioCaps (English)[59]
+Traditional multitask Datasets
+In addition to single-task datasets, there are also comprehensive datasets designed to address multiple speech-understanding tasks.
+These datasets incorporate audio metadata besides the audio-text pair alone.
+For example, the AMI dataset provides rich metadata, including speaker information and discussion topics, making it applicable to a range of tasks such as ASR, Automatic Audio Captioning (AAC), ST, and SD[60].
+Despite their versatility and informational depth, the relatively small size of these datasets limits their broader applicability for training Speech LLMs.
+6.1.2 Datasets Annotated with LLM Participation
+The rapid advancement of Large Language Models (LLM), particularly in text-based reasoning capabilities approaching human performance, has inspired researchers to explore their potential in dataset annotation.
+This approach typically involves representing audio information in text form and pairing it with task-specific prompts to guide LLM in generating inferences.
+This process, known as the SQA format (Speech [in text] + Question + Answer), allows for scalable dataset creation by reducing human labor and lowering costs.
+Furthermore, it provides a novel way to train LLM, facilitating general question-answering capabilities about speech that cannot be fully realized using audio-text pairs alone.
+During training, the text-based question and its corresponding speech context are provided as input, while the model is trained to generate answers in text form as illustrated in 5(a)
+However, despite its efficiency, this method is limited by the reasoning capabilities of the LLM, which may lead to suboptimal annotations in many cases.
+The text prompts and answers generated by LLM often lack the diversity seen in human-created instructions.
+Consequently, such datasets may still fall short in supporting comprehensive speech question-answering tasks.
+6.2 Datasets Generated Using TTS Technology
+Given the scarcity of audio datasets specifically designed for Speech LLMs and their limited variety, researchers have proposed an innovative approach to integrate audio and text modalities more deeply to enable the model’s capability of handling flexible mixed-modal input.
+Building on the SQA framework, this approach converts part of the textual components (speech/questions) into speech using Text-to-Speech (TTS) technology, thereby creating Audio-Text pairs for training Speech LLMs.
+The structure is illustrated in 5(b).
+Peng et al.
+introduced VoiceTextBlender, marking the first significant effort to enhance mixed-modal supervised fine-tuning (SFT) data using TTS[61].
+This method allows prompts and questions to also exist in the audio modality, further enhancing the potential for multi-modal task training.
+This approach opens new possibilities for creating datasets that support deeper cross-modal integration while addressing the limitations of real-recorded audio datasets.
 7 Performance in Speech Tasks
-With the evolution of Speech LLMs, the integration of large language models (LLMs) into speech-based systems has yielded substantial advancements across various dimensions.
-This section investigates the current performance of Speech LLMs in several pivotal tasks in speech understanding, evaluating how LLMs enhance the capabilities of speech models compared to traditional methods.
+With the evolution of Speech LLMs, the integration of large language models (LLM) into speech-based systems has yielded substantial advancements across various dimensions.
+This section investigates the current performance of Speech LLMs in several pivotal tasks in speech understanding, evaluating how LLM enhance the capabilities of speech models compared to traditional methods.
 We will primarily focus on two classic speech-related tasks: Automatic Speech Recognition(ASR), and Speech Translation(ST), each epitomizing a crucial aspect of speech applications.
-Meanwhile, we will further explore the multi-tasking and cross-tasking capabilities exhibited by Speech LLMs across various tasks in the speech understanding field.
+Meanwhile, we will further explore the multitasking and cross-tasking capabilities exhibited by Speech LLMs across various tasks in the speech understanding.
 7.1 Automatic Speech Recognition
-In this survey, we assess the performance of various models via the Word Error Rate (WER) tested on the clean and other test set of LibriSpeech dataset [46], which is an essential benchmark for evaluating ASR systems.
+In this survey, we assess the performance of various models via the Word Error Rate (WER) in English tested on the clean and other test set of LibriSpeech dataset [51], which is a widely recognized benchmark for Automatic Speech Recognition (ASR) systems.
+We also compare two leading models on a multilingual ASR task using the FLEURS test sets [62], another important benchmark for evaluating the multilingual capabilities of ASR models.
 Stage Models
@@
 Whisper Large-V3 1.8 3.6
-Current Speech LLM Stage Qwen-audio 2.0 4.2
-SALMONN 2.1 4.9
+Current Speech LLM Stage Firedasr-LLM 1.73 3.67
+SLAM-LLM 1.84 3.39
 Seed-ASR 1.6 2.8
@@
 For instance, the LSTM-based model Deep Speech 2 achieved a Word Error Rate (WER) of approximately 3.51% on the LibriSpeech clean test set and 10.62% on the other test set [63].
-The introduction of Large Language Models (LLMs) has significantly improved these results.
+The introduction of Large Language Models (LLM) has significantly improved these results.
 Notably, the Conformer model reduced the WER to 1.9% and 3.9% on the clean and other test sets, respectively [25].
@@
 As of 2023, further advancements were seen with Whisper Large-V3, achieving a new low WER of 1.8% on the clean test set.
-Recent Speech LLM models like Qwen-audio [17] and SALMONN [16] have also shown impressive results, approaching state-of-the-art performance, with WERs around 2% on the clean test set and below 5% on the other test set.
+Recent Speech LLMs models like FireRedAsr-LLEM [10] and SLAM-LLM [33] have also shown impressive results, approaching state-of-the-art performance, with WERs around 2% on the clean test set and below 5% on the other test set.
 Currently, the most advanced ASR system on the LibriSpeech dataset is Seed-ASR, which achieves a WER of 1.6% on the clean test set and 2.8% on the other test set [15].
+To evaluate ASR performance across multiple languages, we refer to results from the Seed-ASR report, using the FLEURS dataset [62].
+Compared to Whisper Large-V2, Seed-ASR (ML) demonstrates consistent improvements across all evaluated languages, including Arabic (AR), English (EN), Korean (KO), Japanese (JA), French (FR), and others.
+As shown in Fig.
+6, Seed-ASR outperforms Whisper Large-V2 by a noticeable margin, particularly in high-error languages such as Korean and Arabic, where the gap in WER is most pronounced.
+This highlights Seed-ASR’s robustness and adaptability in multilingual ASR tasks, further solidifying its position as one of the most advanced ASR systems to date.
 7.2 Speech Translation
@@
 In the realm of natural language processing, translation tasks, encompassing both speech translation (ST) and machine translation (MT), are pivotal for enhancing global communication.
-Just like other NLP fields, these translation tasks have experienced significant progress due to the recent advancements in large language models (LLMs) [47].
+Just like other NLP fields, these translation tasks have experienced significant progress due to the recent advancements in large language models (LLM) [64].
 In this survey, we employ the BLEU (Bilingual Evaluation Understudy) score to evaluate the translation performance of each model on the two-way translation between Chinese and English.
@@
 There are two primary objectives in the translation domain: enhancing translation quality and expanding language coverage.
-In machine translation, the NLLB model, which is fine-tuned on LLMs aiming to be a universal translation system, initially broadened its linguistic scope to include more than 200 languages [50].
-Furthermore, BigTranslate, leveraging instruction tuning based on LLaMA, has expanded the translation applicability of LLMs to a broader array of languages, achieving results comparable to those of ChatGPT and Google Translate [27].
-The advent of LLMs has also propelled advancements in speech translation.
+In machine translation, the NLLB model, which is fine-tuned on LLM aiming to be a universal translation system, initially broadened its linguistic scope to include more than 200 languages [66].
+Furthermore, BigTranslate, leveraging instruction tuning based on LLaMA, has expanded the translation applicability of LLM to a broader array of languages, achieving results comparable to those of ChatGPT and Google Translate [13].
+The advent of LLM has also propelled advancements in speech translation.
 Using a Cascaded ASR + MT approach, Whisper has shown notable efficacy in speech translation by utilizing extensive web-scale data [18].
@@
 Please refer to  Figure 7 for a detailed performance comparison of these models.
-6.3 Multi-tasking and Cross-tasking Abilities
-Besides the above-mentioned two speech tasks, some Speech LLMs also demonstrate remarkable results on a wide range of speech tasks, showcasing their good multi-tasking and cross-tasking capabilities.
+7.3 Multitasking and Cross-tasking Abilities
+Besides the above-mentioned two speech tasks, some Speech LLMs also demonstrate remarkable results on a wide range of speech tasks, showcasing their good multitasking and cross-tasking capabilities.
 Among these, Pengi is notable for its versatility, as it supports both open-ended tasks like Audio Captioning and Audio Question Answering (AQA), and close-ended tasks such as Sound Event and Scene Classification, without requiring any task-specific fine-tuning or extensions.
@@
 This model excels in areas like Acoustic Scene Classification, Speech Emotion Recognition, and Automatic Audio Captioning.
-By employing a multi-task training framework that minimizes interference from the diversity of tasks and audio types, Qwen-Audio achieves groundbreaking results in these areas, indicating a significant enhancement in audio understanding capabilities [17].
+By employing a multitask training framework that minimizes interference from the diversity of tasks and audio types, Qwen-Audio achieves groundbreaking results in these areas.
+[19].
+{CJK}UTF8gbsn Input: 甚至出现交易停滞的情况 (A woman says in Chinese, “Even there was almost no transaction” )
+Output
+Ability Test Prompt FireRedLLM Qwen2-Audio-Instruct
+ASR
+Transcribe the audio into text:
+{CJK}UTF8gbsn 甚至出现交易停滞的情况
+{CJK}UTF8minThe original content of this audio is:
+{CJK}UTF8min’甚至出现交易停滞的情况.’
+Speech Translate
+Translate the speech into English:
+{CJK}UTF8gbsn 甚至出现交易停滞的情况
+{CJK}UTF8minThe market even几乎是停滞不前的交易情况。
+Speech Translate
+Translate the speech into Japanese:
+{CJK}UTF8min The speech translated to Japanese is:
+{CJK}UTF8min’甚至出现交易停滞的情况.’
+Speaker Verify
+Is the speaker a female?
+Yes, the speaker is female
+Dialogue Ability
+Based on the speech, can you tell me
+the trading situation?
+The trading situation was almost stagnant.
+Language Detection
+What language is this voice above?
+Mandarin
+Emotion Recognition
+What is the speaker’s emotion?
+The speaker sounds frustrated.
+Speaker Count
+How many speakers are there?
+One
+Dog Detection
+Is there a dog barking in the audio?
+No, there is no dog barking in the audio.
+Speaker Distance
+How far is the speaker?
+The speaker is 25 miles away.
+Speaker Intent
+What is the speaker’s intention?
+The speaker intends to convey that there was almost a
+complete stop in transactions.
+Story Continuation
+Continue the story
+This situation is indeed worrying.
+In the real estate market, a near standstill
+in trading may lead to…
+Table 2: Prompt Tests conducted on FireRedLLM and Qwen2-Audio-Instruct; Text highlighted in gray indicates incorrect outputs, while text in black represents reasonable or correct responses.
 8 Challenges
-Despite significant progress in applying large language models (LLMs) to the field of speech recognition, especially in fixed tasks such as Automatic Speech Recognition (ASR), or leveraging the powerful capabilities of LLMs to accomplish some multitasking, numerous challenges and issues remain.
+Despite significant progress in applying LLM to the field of speech recognition, especially in fixed tasks such as ASR, or leveraging the powerful capabilities of LLM to accomplish some multitasking, numerous challenges and issues remain.
 We validated and tested some of the most cutting-edge current works, which helped to corroborate the problems we identified.
+Fig. 8: The attention score distribution plot for the prompt ”Transcribe speech to Japanese”.
 8.1 LLM’s Dormancy
-Regarding the method of using a projector to project audio feature embeddings into the text token space and combining them with text token embeddings of prompts, although it shows high accuracy and considerable generalization ability in ASR tasks, the LLM does not respond positively to prompts that were not seen during training.
+Regarding the method of using a projector to project audio feature embeddings into the text token space and combining them with text token embeddings of prompts, although it shows high accuracy and considerable generalization ability in ASR tasks, the LLM does not respond positively to prompts that were not seen during training and lost their original intelligent capabilities.
 We refer to this phenomenon as LLM dormancy.
-To further illustrate the issue, we use SLAM-ASR and SALMONN as replication examples:
-The core idea of SLAM-ASR is projecting audio feature embeddings to text token embeddings through a projector composed of a linear neural network.
-Here, we use a recipe consisting of wavlm-large, a projector, and vicuna-7b, with whisper-large and vicuna-7b kept frozen during training while only the projector is trainable.
-We trained for three epochs on the librispeech-960h dataset and achieved results on the librispeech test_clean and test_other datasets close to the expectations outlined in the original paper.
-The core idea of SALMONN is to use a window-level Q-former to integrate audio information extracted by BERT and Whisper and map it into the text token space.
-During training, we updated the parameters of the LLM and encoder using LoRA and the window-level Q-former.
-We achieved results on librispeech test_clean, test_other, and Gigaspeech [52] that matched the expectations of the original paper.
+To further illustrate the issue, we use FireRedLLM[10] and Qwen2-Audio-Instruct[12] as replication examples:
+FireRedLLM is a speech LLM model focused on ASR tasks and has achieved new SOTA results across various recognition datasets.
+The model begins by training a Conformer architecture based on an AED framework, which serves as the audio encoder.
+The encoded speech tokens are then projected into the text embedding space of LLM via an projector module.
+The LLM is fine-tuned using LoRA to adapt its capabilities for the specific task.
+Notably, FireRedLLM is exclusively trained for ASR purposes, with the training prompts remaining fixed throughout the process.
+Qwen2-Audio-Instruct is a multitask speech model capable of handling speech understanding tasks like speech recognition and translation, as well as conversational tasks.
+It uses the encoder from Whisper as its speech encoder and projects the encoded tokens into the text embedding space of LLM via a projector layer.
+Training is done in three stages: first, ASR and AAC tasks are trained; second, instruction fine-tuning is performed using various speech understanding tasks; and third, RLHF fine-tuning is applied to align with human conversational preferences.
+Notably, the training data for Qwen2-Audio-Instruct amounts to 600,000 hours.
 We conducted a series of prompt tests on the two models mentioned above, with the results as shown in Table 2.
-From the table, we can see that SLAM-ASR almost does not respond to prompts, while SALMONN selectively responds to prompts, which precisely confirms our view that LLMs cannot maintain full model performance in the current framework.
-Input: Can you divorce life from art?(A man speaker with chinese accent)
-Output
-Ability Test Prompt SLAM-LLM SALMONN
-ASR transcribe speech to english text Can you divorce life from art? can you divorce life from art?
-Speech Translate transcribe the speech to text and translate it to chinese. {CJK}UTF8gbsn你能从艺术中离开吗?
-Speech Translate transcribe the speech to text and translate it to Japanese. {CJK}UTF8minアートから生活を破棄することができますか?
-Speaker Verify which scene will this talk occur? The speaker is male and has an American accent. He speaks with confidence and clarity.
-Speaker Verify Is the speaker a female? Yes, the speaker is a female.
-Dialogue Ability Tell me the reason. I’m sorry, I’m not sure what you’re asking. Could you please provide more context or clarify your question?
-Dialogue Ability
-Based on the speech, can you help me determine
-the answer?
-Yes, I can divorce life from art.
-Dialogue Ability can you answer the speech and give reasons? Can you divorce life from art
-Context Inference
-Actually this is a speech a man asked me. I don’t
-have answer. And someone tells me to say yes.
-But I don’t know why. Can you explain it?
-Can you divorce life from art?
-Yes, I can explain it. The man asked you if you could
-divorce life from art. This means that you can separate
-the concept of life from the artistic representation of it.
-For example, a painting of a landscape could be seen
-as a representation of life, but not necessarily a reflection
-of it. So, the answer to the man’s question is yes, you can
-divorce life from art.
-Table 2: Prompt Tests conducted on Slam-ASR and SALMONN; Text highlighted in red indicates errors or incorrect outputs, while text in black represents reasonable or correct responses.
-From the perspective of training strategies, SLAM-ASR did not update parameters for the encoder or LLM, thus making it more intuitive that the problem likely arises from the intermediate modality fusion step.
-To validate our approach, we attempted to reproduce the training process of SLAM-ASR.
-We then compared the average absolute value of the projected audio feature vectors with the average absolute value of the token embeddings obtained from encoding the prompt.
-Fig. 7: Mean absolute value of embeddings
-Note: As the tensor shape of the input embedding is (Batch size, Sequence id, Feature dimension), we set batch size to 1 and calculate the average of the absolute value of all the elements corresponding to the sequence id.
-The figure here is a randomly chosen audio input for clear demonstration.
-From Figure 7, we observe that the audio embeddings are significantly larger compared to the text embeddings.
-Referring back to the procedure in the transformer module in a typical large language model, the text is first converted into tokens and embeddings before calculating the attention mechanism, which includes positional embeddings.
-However, the audio feature vectors are projected and then directly concatenated with the text token embeddings in Speech LLM.
-Clearly, there is no attempt to capture the positional information of the two modalities here.
-A second potential issue arises during the calculation of attention scores.
-Recalling the standard steps involved in calculating attention scores in the attention mechanism, Given the audio-text modality embedding vector $E$, and since the LLM is kept frozen, the weight matrices $W^{Q}$, $W^{K}$, and $W^{V}$ are fixed.
-We transform $E$ into query, key, and value vectors using linear transformations, calculated as follows:
-$Q$ $=EW^{Q}$ (3)
-$K$ $=EW^{K}$ (4)
-$V$ $=EW^{V}$ (5)
-Next, we compute the attention scores:
-Attention $\text{Attention Scores}=QK^{T}$ (6)
-where $K^{T}$ is the transpose of the key vector.
-To stabilize the training and inference process, we optionally scale the dot-product attention scores:
-Scaled $\text{Scaled Scores}=\frac{QK^{T}}{\sqrt{d_{k}}}$ (7)
-where $d_{k}$ is the dimensionality of the key vectors.
-The scaled attention scores are then transformed into probability distributions using the softmax function:
-[…]
-Finally, we use the attention weights to perform a weighted sum of the value vectors to obtain the final attention output:
-Attention Output=Attention $\text{Attention Output}=\text{Attention Weights}\cdot V$
-which in detail can be expressed as:
-Attention $\text{Attention Output}=\left(\text{Softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)\right)\cdot V$ (9)
-Here we can observe that since the large language model is frozen, only $E$ changes based on the input.
-The embedding $E$ is obtained by concatenating the audio feature embedding $A$ and the text feature embedding $T$:
-$E=[A\;T]$
-We can also decompose $Q$, $V$, and $K$ into vectors corresponding to $A$ and $T$, for example,
-$Q=[Q^{A}\;Q^{T}]$
-$V=[V^{A}\;V^{T}]$
-$K=[K^{A}\;K^{T}]$
-Since $A\gg T$, according to equations (1), (2), and (3), due to the shared weight matrices, $Q^{A}\gg Q^{T}$ and similarly for other vectors.
-It is clear that the final attention output from the speech part is much larger than that from the text part.
-As a result, the model is likely to neglect the text portion, leading to the large language model falling into dormancy.
-This can also be explained more broadly and intuitively using Bayes’ theorem [16]. Let $Y$ denote the output, $X$ denote the audio feature, and $I$ denote the instruction.
-$P(Y|X,I)=\frac{P(Y|X)\cdot P(I|Y,X)}{P(I|X)}$ (10)
-If we provide only one or a few corresponding outputs $Y$ for a given audio input, then:
-$\frac{P(Y|X,I)}{P(Y|X)}$ (11)
-is not negligible. In other words, the influence of $I$ on the result will not be significant.
-Models like SALMONN and Qwen-audio, which are Speech LLMs, perform poorly in tasks involving untrained speech emotion inference and environmental understanding [53], and may even provide irrelevant answers.
+From the table, we can see that FireRedLLM almost does not respond to prompts, while Qwen2-Audio-Instruct incorrectly answered some questions, which precisely confirms our view that LLM cannot maintain full model performance in the current framework.
+Why do speech large models lose their original intelligence?
+As figure 8 shows, this can be seen from the attention scores assigned to speech tokens versus text tokens.
+The FireRedASR model, trained solely on ASR tasks, almost entirely focuses on speech tokens.
+In contrast, the Qwen2-Audio-Instruct model, which uses multi-task training, shows some improvement but still exhibits a stronger preference for speech tokens compared to text tokens.
+This can also be explained more broadly and intuitively using Bayes’ theorem [20].
+Let $Y$ denote the output, $X$ denote the audio feature, and $I$ denote the instruction.
+Instruction I is either completely fixed or selected from a limited set of tasks.
+Compared to the input audio content, instruction I lacks diversity.
+$P(Y|X,I)=\frac{P(Y|X)\cdot P(I|Y,X)}{P(I|X)}\approx P(Y|X)$ (3)
+Other models like SALMONN and Qwen-audio also perform poorly in tasks involving untrained speech emotion inference and environmental understanding [68], and may even provide irrelevant answers.
+SALMONN refers to this phenomenon as task-overfit.
+SALMONN’s research reveals that decreasing the coefficient of Lora can enhance the diversity of model responses and mitigate task-overfit.
+However, this approach may lead to a decline in training task outcomes.
+To address this, a new activation fine-tuning stage with few shot samples was introduced.
+The results indicate that this method can reduce the occurrence of model task-overfit, yet the problem has not been completely resolved.
 This further supports the observation that we did not fully achieve modality alignment under the current framework.
@@
 Considering the diversity of inputs and outputs in large language models, exploring the incorporation of reinforcement learning strategies such as Proximal Policy Optimization (PPO) [49] can be beneficial.
-Introducing perturbations in supervised learning can help maintain the model’s multi-task performance.
+Introducing perturbations in supervised learning can help maintain the model’s multitask performance.
 Additionally, to address hallucinations in large language models, techniques such as Reinforcement Learning with Human Feedback (RLHF) can be employed to apply ”soft” constraints [69], alleviating issues such as repetitive outputs and other hallucination phenomena.
@@
 10 Conclusion
-This work presents a comprehensive exploration of Speech Large Language Models (Speech LLMs), focusing on their architecture, training strategies, and advancements in the field of Spoken Language Understanding (SLU).
+This work presents a comprehensive exploration of Speech Large Language Models (Speech LLMs), focusing on their architecture, training strategies, and advancements in the field of Speech Understanding.
 We examined the evolution from traditional methods to novel architectures that integrate Audio Feature Extraction, Multimodal Information Fusion, and LLM Inference, facilitating richer audio processing and deeper multimodal understanding.
@@
 The survey highlights key improvements in Rich Audio Transcription and the potential for Cross-task Integration across SLU tasks.
-However, challenges such as the Dormancy of LLMs under certain conditions persist, underscoring the need for further innovation in training strategies and system design.
+However, challenges such as the Dormancy of LLM under certain conditions persist, underscoring the need for further innovation in training strategies and system design.
 By addressing these challenges and exploring future solutions, this work provides valuable insights into advancing Speech LLMs and their applications in multimodal contexts.